### v2

# LLMs Capture Urban Science but Oversimplify Complexity

Abstract
Large language models (LLMs) are increasingly used in scientific reasoning and data generation tasks, yet their alignment with empirical knowledge and practical urban data in urban science remains unclear. Here, this study introduces AI4US (Artificial Intelligence for Urban Science), a framework for systematically evaluating five leading LLMs (ChatGPT, DeepSeek, Claude, Gemini, and ChatGLM) by generating urban data and testing their fidelity against real data distributions and the empirical results of three foundational theories, including urban scaling laws, urban distance decay, and urban vitality. LLMs reproduce core theoretical patterns with notable accuracy (average $\mathrm{R}^{2}=0.804$ for scaling; $0.988$ for decay), and exhibit capacity to simulate less formalized constructs such as urban vitality. However, generated data exhibit crucial limitations: poor diversity compared to real data; systematic deviations in fitted theoretical parameters (e.g., underestimated scaling exponents and decay rates); and limited improvement from prompt engineering. These findings suggest that while LLMs offer a promising research paradigm for exploring urban science, they fall short of capturing real-world complexity, underscoring the need for improved alignment and evaluation.
Machine Learning, ICML
1 Introduction
Urban science has long sought to decipher the complex behaviors of cities from early demographic surveys to computational modeling [1, 2, 3, 4]. In recent decades, an explosion of urban data from remote sensing, sensor networks, and digital platforms has promised to revolutionize this endeavor [5, 6]. Yet, a fundamental paradox remains. Despite this data abundance, theoretical innovation has stagnated [7, 8, 9]. The field struggles to move beyond fragmented, context-dependent insights toward broadly generalizable urban theory. This impasse stems from a persistent methodological gap. The field lacks scalable and systematic experimental designs that allow for robust, cross-city replication and validation of scientific claims [8, 10, 11].
This theoretical stagnation is rooted in the inherent difficulty of conducting controlled experiments in urban settings. Historically, urban experimentation has evolved through several paradigms, each with significant limitations. Early approaches treated the city as a living laboratory, relying on direct observation and policy trials, but these methods lacked scalability and control [12, 13]. Physical and analogue models offered controllable surrogates but were often oversimplified, failing to capture the non-linear dynamics of human behavior at a meaningful scale [14, 15]. The subsequent rise of digital models, such as cellular automata and agent-based simulations, introduced rule-based experimentation but hinged on hard-coded assumptions and static calibration, limiting their ability to represent the nuanced, adaptive behaviors of individuals [16, 17, 18]. These shortcomings are magnified by the profound transformations of contemporary cities, from autonomous mobility to platform economies, creating an urgent need for a new experimental paradigm capable of capturing the complex human-in-the-loop dynamics that define modern urban life [5, 19].
Nowadays, advances in generative AI are spearheading a new paradigm of data-driven scientific discovery [20, 21, 22]. In fields like biology and materials science, its success lies in predicting specific physical states from vast solution spaces, exemplified by AlphaFold’s protein structure prediction [23] and MatterGen’s discovery of novel materials [24]. The challenge in urban science is fundamentally different. It is not to predict a discrete physical outcome [25], but to capture the broad statistical patterns that emerge from complex, adaptive urban systems, which are governed by context-dependent social behaviors rather than universal physical laws [26]. This distinction positions LLMs as a uniquely compelling tool. Trained on diverse records of human behavior and expression, LLMs offer a powerful pathway to generate synthetic urban data with high fidelity, enabling scalable urban experimentation where formal modeling has reached its limits.
In urban science, however, the potential of generative AI to function as a new form of virtual city laboratory that can produce experimental data [27, 28], analyze urban phenomena, and drive theoretical advancement remains largely unexplored [20, 29]. Initial studies have shown promise, using LLMs to generate synthetic mobility data, simulate residents’ daily behaviors, or assess urban systems [22, 30]. However, these pioneering efforts often remain fragmented. A significant research gap persists in systematically validating the fidelity of these worlds generated by AI against established urban theories [31]. Furthermore, the inherent geographical and social biases within LLMs pose a critical challenge to their reliability as scientific instruments [32, 33]. Therefore, a structured and rigorous framework is necessary to move beyond isolated applications and systematically evaluate the potential and pitfalls of LLMs as a foundation for the next generation of urban science.
To address these gaps, we introduce the AI4US (Artificial Intelligence for Urban Science) framework. World models are internal, generative representations that simulate real-world dynamics [33]. This framework pioneers a methodology to evaluate LLMs as a first-generation proxy for urban world models. Our approach probes the implicit urban knowledge embedded within LLMs by testing their ability to reproduce three foundational urban theories: urban scaling laws, distance decay relationships, and Jane Jacobs’ urban vitality theory [29, 34, 35]. These theories were chosen to create a comprehensive testbed, as they represent fundamental urban mechanisms across different scales and dimensions: urban scaling laws govern macro level regularities, distance decay explains meso scale spatial interactions, and urban vitality captures micro scale human environment dynamics [36, 37, 38, 39, 40, 41]. Our findings reveal that while LLMs can replicate these theoretical patterns with notable fidelity, they also exhibit systematic limitations, underscoring that the path toward true theoretical discovery will require a new generation of world models capable of deeper causal reasoning.
2 Results
2.1 LLMs capture the fundamental forms of urban theories
Figure 1: Urban scaling law reproduced on generated data with LLMs: generated versus empirical coefficients. “Replicate Tests” contains 100 independent duplicate data generation experiments for each LLM, and each experiment generates data of 100 cities, among which GPT-4o has the best performance (see Supplementary Experiment 6). The diagram shows the results of GPT-4o.
Figure 2: Urban decay reproduced on generated data with LLMs: generated versus empirical statistical and spatial patterns. a, Concentric-ring land-density profiles for three representative cities on different continents across three decades (1990–2010), together with the corresponding “Mirage Cities” produced by GPT-4o when both city name and time period are included in the prompt. b, Sensitivity of the generated decay curves to prompts that specify different maximum radii and historical periods. c, Direct side-by-side comparison of Beijing’s empirical profile with that of its Mirage City generated by LLMs.
Figure 3: Urban vitality revealed on generated data with LLMs. “Replicate Tests” contains 100 independent duplicate data generation experiments for each LLM, and each experiment generates data of 100 blocks. The diagram shows the results of GPT-4o.
Figure 4: Data distribution divergence analysis. (a) Equal-width binning with relative frequency normalization was applied to convert raw sample counts into scale-invariant frequency distributions. Mean Absolute Error (MAE) of bin-level frequencies quantified the central tendency divergence between real data (R) and generated data (G), calculated as the sum of absolute differences across all 15 bins. Overlap Ratio (OR) measured quantile interval congruence between corresponding bins of R and G distributions. (b) Jensen-Shannon Divergence (JSD) comparisons among four datasets: real data (R), baseline generated data (G), China-prompted generated data (GC), and USA-prompted generated data (GA). The inclusion of geographical constraints, such as “China”, increased GPT-4o’s constraint-triggering rate by 24% (24/100 trials), e.g., “I can’t provide real-time data such as the current population”.
Figure 5: AI4US, the proposed framework utilizing LLMs to advance urban science.
The investigation reveals that LLMs possess a remarkable ability to recognize and reproduce the foundational forms of established urban theories across multiple scales, from macro-level city systems to micro-level neighborhood dynamics. At the macro-scale, when tasked with generating data for urban scaling laws, LLMs consistently produce outputs that adhere to the characteristic power-law relationship. The generated data exhibits a strong fit to the scaling law equation, achieving a high average coefficient of determination ($R^{2}=0.804$) (Fig. 1a). Most models successfully replicated the expected super-linear relationship ($\beta>1$) for socioeconomic output and, in the case of GPT-4o, the sub-linear relationship for infrastructure. Moving to the meso-scale, the models display a non-trivial grasp of the concentric-zone structure posited by distance-decay theory. When generating land density surfaces, the outputs fit the inverse-S formulation with exceptional accuracy, showing a mean $R^{2}$ of 0.988 (Fig. 2a).
Furthermore, this capability extends beyond quantitative laws to less-formalized, qualitative theories. At the micro-scale, when confronted with urban vitality theory, LLMs demonstrate a clear alignment with its core tenets. By generating data for key indicators and then assessing their correlation with a livability score, the models produced relationships directionally consistent with Jane Jacobs’ theory (Fig. 3a). Specifically, indicators such as ’Building Mix Index’ and ’Tall Building’ showed the highest consistency with the theory, while ’Building Mix Index’ and ’Population Density’ demonstrated the most significant positive impact on the generated livability scores. In summary, LLMs can consistently generate data that is structurally ”form-similar” to a diverse set of urban theories, whether they are mathematically explicit or qualitative, establishing a baseline of theoretical competence.
2.2 Mirage cities systematically oversimplify real-world complexity
While LLMs can replicate theoretical forms, a deeper analysis reveals a significant gap between their generated ”Mirage cities” and the complexity of real urban systems. This discrepancy manifests as a series of systematic, non-random deviations. A primary finding is the lack of diversity in the generated data compared to empirical data. The analysis of data distributions reveals that LLM-generated data is significantly more homogeneous. Without specific geographic prompts, the generated data’s numerical scale diverges substantially from real data, resulting in a low average Overlap Ratio (OR) of just 28.6% per bin (Fig. 4a). More critically, the Jensen-Shannon Divergence (JSD) between real-world data samples (R-R) is, on average, 122.54% higher than the divergence within generated samples (G-G) (Fig. 4b). This indicates that real data exhibits substantially greater internal variation. This inherent ”over-smoothing” tendency is a foundational reason why LLM-generated data, while theoretically clean, lacks the real-world noise and heterogeneity essential for robust scientific analysis.
This lack of diversity contributes to specific parametric and structural distortions when fitting theories. At the macro-scale, a systematic underestimation of the scaling exponent ($\beta$) is observed in the urban scaling law test (Fig. 1b). The models consistently produce lower exponents for both socioeconomic output and infrastructure compared to established empirical values (e.g., $\beta\approx 1.15$ for GDP and $\beta\approx 0.85$ for infrastructure), suggesting they capture the ”shape” of the law but fail to grasp its true ”steepness” or magnitude. At the meso-scale, this manifests as the generation of ”idealized” spatial decay curves that are smoother than reality (Fig. 2a, Fig. 3b). The generated data gravitates toward the theoretical extrema, with values approaching 1 at the center and 0 on the periphery. Consequently, this idealized output fails to capture the ”messiness” of real cities, which contain features like central parks or suburban industrial belts that cause deviations from a perfect curve. The distributional difference in these generated decay curves is 149% smaller than that of empirical data, further underscoring their idealized nature.
2.3 The potential for improving AI-generated data
Given the observed deviations, this study investigated whether these gaps could be bridged through different forms of intervention. The results indicate that the fidelity of generated data possesses potential for improvement through both user interaction and model selection. Strategic prompting offers a viable pathway to enhance data fidelity. Prompting with specific geographic contexts demonstrates a clear effect on data distribution; when prompted with ”China,” the generated data showed a greater alignment with real data from Chinese cities, increasing the Overlap Ratio by 38.56% and decreasing the Mean Absolute Error by 52.14% (Fig. 4a, 4b). Similarly, at the meso-scale, prompting with parameters like city size or historical period directly influences the resulting decay curves (Fig. 2c). These results confirm that the model’s output is not static but sensitive to contextual cues, highlighting a clear potential for iterative refinement.
Furthermore, performance varies noticeably across the different LLMs evaluated. In the urban scaling law test, GPT-4o demonstrated the closest alignment with empirical results, while Claude 3.5 was most accurate in reproducing Zipf’s law. For distance decay, GPT-4o again showed more nuanced spatial reasoning, whereas models like ChatGLM-4 and DeepSeek-V3 tended to produce less diverse outputs. In the urban vitality assessment, GPT-4o consistently outperformed all other models. These variations suggest that factors intrinsic to the models, such as architecture, training data, and fine-tuning strategies, significantly influence the characteristics and fidelity of the generated urban data.
2.4 A new paradigm for qualitative theory exploration
The test of urban vitality, a theory that is descriptive and lacks a single mathematical form, reveals a distinct capability of LLMs, shifting their role from a ”physical simulator” to a ”socio-cultural simulator.” By prompting an LLM to act as a ”resident” evaluating a neighborhood, the task is not to solve a physical equation, but to have the model tap into its vast training data of human text and discourse to simulate a ”social consensus” on what makes a place livable. The model’s outputs revealed nuanced and theoretically consistent insights (Fig. 3). For instance, it identified ’Building Mix Index’ and ’Tall Building’ as having the highest consistency with vitality theory, while ’Aged Building’ showed the greatest deviation. In terms of impact, it judged ’Building Mix Index’ and ’Population Density’ as having the most significant positive effect on livability scores (Fig. 3b). This demonstrates that LLMs can serve as a powerful ”hypothesis generation engine.” They can process the complex, multi-faceted information embedded in qualitative theories and produce structured, testable hypotheses about the relative importance of different urban elements. This capability to ”operationalize” qualitative knowledge marks a significant new potential for AI in advancing urban science.
3 Discussion
This study reveals a fundamental duality in the capabilities of current Large Language Models within urban science. They excel at reproducing the abstract, mathematical forms of established theories but falter in capturing the statistical complexity and fidelity of real world urban systems. The systematic evaluation across different scales demonstrates that while LLMs can generate data that is structurally ”form similar” to urban theories, this apparent competence masks significant deficiencies in diversity and parametric accuracy. These findings position LLMs not as high fidelity physical simulators, but as powerful engines for exploring the socio-cultural patterns embedded in human language, a role that offers both profound opportunities and critical limitations for scientific inquiry.
The initial finding that LLMs consistently generate ”mirage cities” with low diversity and systematically distorted parameters (e.g., underestimated scaling exponents) points to a core limitation. This phenomenon, where models produce archetypal representations of cities, a tendency sometimes described as ”group flattening” [31], aligns with observations of geographical and social bias in other domains [32, 33]. This interpretation suggests that these distortions are not random errors but are symptomatic of the LLMs’ nature as statistical pattern matchers rather than causal world models. Lacking a deep, generative understanding of the mechanisms that produce urban phenomena, they instead reproduce the most dominant statistical patterns from their training data, smoothing over the granular heterogeneity and outliers that are often crucial drivers of real world dynamics. This tendency, also noted in other applications of LLMs for generating synthetic populations [31], poses a significant risk for empirical research that relies on capturing the full spectrum of urban complexity.
The investigation into data fidelity reveals that LLMs are not static instruments but are malleable tools sensitive to user interaction and model specific characteristics. Our finding that prompt engineering can partially mitigate deviations by steering the model toward a specific context underscores the importance of the human AI interface in scientific applications [30, 31]. The quest is not for a single, perfect prompt, but for a methodology of iterative refinement where the prompt itself is a key parameter in the experimental design [42]. Furthermore, the observed performance differences across the five evaluated LLMs highlight that model architecture and training strategies are critical variables. This suggests that future improvements in data fidelity will depend on both more sophisticated interaction protocols and the rapid advancements in the models themselves [20].
The limitations of LLMs also unveil a novel scientific opportunity. Their tendency to generate theoretically ”clean” data, stripped of real world noise, makes them powerful instruments for theoretical exploration [29, 34, 35, 31, 33]. Urban scientists can now generate vast, fine grained datasets under specific theoretical assumptions, creating virtual laboratories to explore a theory’s logical consequences or test its sensitivity to certain parameters. This moves beyond simply validating models against existing data and toward a new mode of inquiry focused on understanding the internal consistency and emergent properties of theories themselves. This approach directly addresses the challenge of how AI can generate new knowledge for urban studies, by providing a scalable method to rigorously interrogate theoretical constructs in a controlled environment.
A distinct capability of LLMs is revealed when they are tasked with navigating less formalized, qualitative theories like urban vitality. Here, the LLM is not solving a physical equation but is prompted to simulate a ”social consensus” by synthesizing the vast amount of information about cities embedded in its training data, from academic texts to public discourse. The result is not a ”discovery” of a new law, but the generation of a plausible, structured set of relationships that can provide urban scientists with novel, testable hypotheses. This aligns with the growing recognition of LLMs as tools for urban sensing and knowledge synthesis. In this role, the LLM acts as a powerful catalyst for human led scientific inquiry, offering a new method to ”operationalize” qualitative knowledge into a quantitative, empirical research program.
While this study provides a foundational framework, its scope defines a clear agenda for future research. Continuous benchmarking is essential, as the quantitative performance of LLMs will evolve with new architectures [20]. More importantly, our findings underscore a critical distinction. The systematic biases and lack of causal depth suggest that current LLMs, while powerful, are not true world models. They are adept at learning the ”what” from data but struggle with the ”why”. The path toward transformative breakthroughs in urban science, particularly in areas like causal inference and policy simulation [43], will likely require a new generation of hybrid models. Future work should focus on integrating the rich, contextual knowledge of LLMs with the causal integrity of traditional urban simulation frameworks, a synergy that could unlock the next frontier of a robust and predictive urban science.
4 Method
The intelligence of large language models essentially derives from their ability to compress world knowledge acquired during training [44, 45]. To systematically evaluate the capabilities of LLMs in urban science, this study employs a framework grounded in three foundational urban theories (Fig. 5). The selection of these theories is deliberate, designed to test the LLMs’ generative fidelity across the distinct yet interconnected scales at which urban phenomena are understood: the macro-level system of cities, the meso-level internal structure of a single city, and the micro-level dynamics of a neighborhood.
Substantial evidence suggests that, similar to regularities observed in natural complex systems, urban phenomena exhibit patterns and order across these multiple dimensions of size, space, and morphology [36, 39, 40, 41]. The chosen theories are canonical representations of these scaled phenomena (Table 1). Urban scaling laws capture nonlinear regularities at the macro-scale, describing how infrastructure and socioeconomic activity grow with city size. Distance decay theory reflects meso-scale spatial interaction friction and geographic reach, explaining the internal organization of cities. Finally, Jane Jacobs’ theory of urban vitality articulates how fine-grained, micro-scale morphological configurations contribute to emergent urban dynamics. Together, these theories exemplify foundational mechanisms in urban systems. They provide a comprehensive and representative testbed for evaluating the core capabilities of LLMs: from reproducing mathematically explicit laws to simulating the socio-cultural knowledge embedded in qualitative theory.
Table 1: A Multi-Scale Framework for Evaluating LLMs in Urban Science.
Scale Theory Significance in Urban Science
Macro Urban scaling law Explains universal, size-based regularities of cities as
(City system) complex systems, central to understanding urban growth
and efficiency.
Meso Urban distance decay Represents the foundational role of spatial friction and
(Internal city structure) accessibility in shaping internal urban structure and
organization.
Micro Urban vitality Provides a human-centric perspective linking the micro-design
(Block) of the built environment to social and economic outcomes;
a cornerstone of urban design.
4.1 Comparison of data distribution
First, this study directly compares the data distributions of generated and real-world samples. Since LLM-generated values without geographic prompts often differ from real samples by orders of magnitude, this study applies equal-width binning and relative frequency normalization to eliminate differences in data scale, converting the raw sample counts within each bin into frequencies. The Mean Absolute Error (MAE) of bin-level frequencies is then used to measure the average difference in relative frequency between real and generated data across bins, serving as an indicator of how well the generated data approximates the overall distribution trend of the real data. This is calculated by summing the absolute differences of mean frequencies across all bins, thereby capturing the degree of deviation in central tendency between the two distributions. For each bin, the average frequency is computed separately for the real data (R) and the generated data (G). $N$ is the number of bins, $\mu_{R,i}$ is the average relative frequency of R in the ith bin, $\mu_{G,i}$ is the average relative frequency of G in the i th bin.
$\text{MAE}=\frac{1}{N}\sum_{i=1}^{N}|\mu_{R,i}-\mu_{G,i}|$ (1)
Measure the overlapping degree of quantile intervals between real data and generated data in each bin, and evaluate whether the fluctuation range of generated data covers the distribution interval of real data. The overlapping ratio calculates the ratio of the interval intersection of the two to the union, and divides the sum of the intersection areas of all bins by the sum of the union areas.
$\text{Overlap Ratio}=\frac{\sum_{i=1}^{N}\text{Intersection}(Q_{R,i},Q_{G,i})}{\sum_{i=1}^{N}\text{Union}(Q_{R,i},Q_{G,i})}$ (2)
Secondly, based on the urban data generated by LLMs and the real data of Chinese cities, this study measures the pairing difference between the two data and three discrete distributions by Jensen-Shannon Divergence (JSD). Include ’distance between real sample and real sample data distribution’ (R-R), ’distance between real sample and generated sample data distribution’ (R-G), ’Distance between generated sample and generated sample data distribution’(G-G). Taking the calculation of average divergence of R-G as an example, there are $n$ true distributions $R_{i}$ and $m$ generated distributions $G_{j}$.
$\text{Distance}_{R-G}$ $=\frac{1}{n\times m}\sum_{i=1}^{n}\sum_{j=1}^{m}\text{JSD}(R_{i},G_{j})$ (3)
$\text{JSD}(R,G)$ $=\frac{1}{2}\text{KL}(R\|\frac{1}{2}(R+G))+\frac{1}{2}\text{KL}(G\|\frac{1}{2}(R+G))$ (4)
$\text{KL}(R\|M)$ $=\sum_{x}R(x)\log\frac{R(x)}{M(x)}$ (5)
4.2 Urban scaling law
Scaling laws are a fundamental concept in complex systems, describing how certain quantities change with system size, typically following a power-law relationship. In the context of urban systems, urban scaling laws serve as functional models for understanding cities as complex systems. According to Ribeiro and Rybski [46], these scaling laws can be categorized into two broad types: inter-urban processes (models describing interactions between cities) and intra-urban processes (models describing dynamics within a single city). Inter-urban processes focus on the exchange of information between cities as a mechanism for explaining the emergence of scaling laws. Key theoretical explanations for inter-urban scaling include Zipf’s Law, hierarchical organization, and interactions between people in different cities. In contrast, intra-urban processes examine scaling laws based solely on factors within a city. These models explore how human interactions, urban geometry, and spatial distribution contribute to urban scaling. Within this category, gravity models represent a generalized formalization of intra-urban scaling dynamics.
Empirical studies on inter-urban processes have shown that city rank (Y) and city population size (N) exhibit a scaling relationship [46]. For intra-urban processes, empirical evidence suggests two distinct scaling laws: (1) variables related to socioeconomic activity (e.g., GDP, patents, HIV cases) exhibit super-linear scaling ($\beta>1$), meaning that as a city’s population increases, these socioeconomic indicators grow at a faster rate; and (2) variables related to infrastructure (e.g., the number of charging stations or gas stations) exhibit sub-linear scaling ($\beta<1$), indicating that as the population increases, these infrastructure-related variables grow at a slower rate.
For Zipf’s Law between cities, the relationship typically follows a power-law function where $\beta\approx-1$, meaning that city size (e.g., population) and city rank follow an inverse relationship. In this equation, $\log_{10}(a)$ represents the intercept, corresponding to the constant $a$ in the power-law function, while the slope $\beta$ determines the exponent of the power-law relationship. Here, $y$ represents population size, and $x$ denotes city rank. The model incorporates several statistical parameters: $n$ refers to the number of independent trials fitting the theoretical equation, $g$ represents the number of generated data samples per trial, and $m$ denotes the parameters of the theoretical equation (such as $a$ and $\beta$ in Zipf’s Law between cities). The standard deviation of $m$ across $n$ trials reflects variability in parameter estimation, while the mean value of $m$ provides an overall estimate. A narrower confidence interval (CI) for $\beta$ suggests a more precise estimate. Additionally, the coefficient of variation (CV) for Zipf’s Law parameters is computed as the geometric mean of the CV for $a$ and $\beta$, assuming that these parameters are independent and without considering their correlations.
$\log_{10}(\text{Population})=\log_{10}a+\beta\log_{10}(\text{Rank})$ (6)
$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (7)
The urban scaling law in cities describes how the infrastructure and socio-economic factors in a city change with the growth of the city scale. It is the scaling law of urban infrastructure and urban social factors derived by Bettencourt in 2013 based on the Human Interaction model. $Y$ represents the infrastructure or socio-economic elements in the city. $\text{CV}_{\text{Urban}}$ represents the coefficient of variation of the parameters of the urban scaling law in cities under $n$ independent experiments.
$\log_{10}(y)=\log_{10}a+\beta\log_{10}(\text{Population})$ (8)
The understanding of Zipf’s Law and urban scaling law by the LLMs was tested for inter-city and intra-city models, respectively. By inputting the appropriate prompt, the large model is used to generate city samples, including city serial number, population, road length, and GDP, and the theoretical equations of inter-city model and urban internal model are fitted respectively, and the goodness of fit and regression coefficient are observed, and the differences between different large models are compared. And compare it with the value of empirical research [38]. Each model repeats the independent experiment 10 times to generate data samples with averaged metrics.
Table 2: Prompt and specific description for urban scaling law.
Prompt Description
Urban scaling law Generate the dataset containing 100 cities in a country. Format the output as: CityName1, Population1, Infrastructure volume1, GDP1 CityName2, Population2, Infrastructure volume2, GDP2 Output exactly 100 lines without any additional text or explanations. - City Name - Population - Infrastructure volume: Total road miles - GDP: all Gross Domestic Product of the city in one year
4.3 Urban distance decay
The term distance decay is used to describe the diminishing intensity of interaction between two locations as the distance between them increases. It refers to the phenomenon where the strength of connections — especially in economic activities — tends to weaken with greater spatial separation [47]. At its core, the distance decay law posits that interactions between geographic entities are inversely related to distance; under otherwise equal conditions, such interactions typically follow an inverse-square relationship with distance. As a concept that reflects the decline of influence over space, distance decay is closely related to Tobler’s First Law of Geography [48]: ’Everything is related to everything else, but near things are more related than distant things’.
In urban geography, cities and towns often have clearly defined centers, and most urban attributes — such as population density, land value, and the quality or quantity of services — tend to decrease with increasing distance from the center [47]. Due to spatial agglomeration, urban spatial structures frequently exhibit a pronounced ’center-periphery’ pattern, where the intensity of urban elements decays with increasing distance from the city core. The most typical manifestation of this is the spatial decay of population density. Common quantitative models include the negative exponential model [49] and the inverse power model [50]. For urban land use density, the inverse S-function model has also been proposed to characterize its spatial decay [51].
For urban land density, the inverse S-function model describes a decaying distribution pattern that follows an inverse S-shaped curve as one moves outward from the city center. In this model, $L$ represents the saturation density parameter, indicating the maximum possible land density (i.e., the asymptote of the curve). $k$ denotes the growth rate parameter (1/km), which reflects how rapidly the density transitions from low values to the saturation level; the larger the value of $k$, the steeper and more abrupt the transition. $x_{0}$ is the center point parameter (km), representing the location at which the land density reaches half of its maximum value — essentially, the inflection points of the curve. $\mathrm{CV}_{\mathrm{DLand}}$ represents the coefficient of variation of the related parameters of land density attenuation in the distance decay theory under $n$ independent experiments.
$f(r)=\frac{1-c}{1+e^{a\left(\frac{2r}{D}-1\right)}}+c$ (9)
The extent to which current LLMs understand the spatial distribution of population density and land density according to the distance decay law was assessed. By providing appropriate prompts, LLMs simulated a mirage city with 100 concentric zones and generated samples that included zone identifiers, population density, and land density attributes for each zone. Then the study fitted theoretical equations for both population and land density to the generated data, evaluating the goodness of fit and regression coefficients, and compared these values with those found in empirical studies [52]. Additionally, the study examined the differences across various LLMs. Each model conducted 100 independent trials to generate data samples, and finally reported the averaged results across these repetitions.
Table 3: Prompt and specific description for urban distance decay theory.
Prompt Description
Urban distance decay theory Generate attributes for 100 concentric city circles arranged in order in a city. From the city center (Circle 1) to the outermost ring (Circle 100). Format the output as: Circle 1, Population Density1, Land Density 1 Circle 2, Population Density2, Land Density 2 Output exactly 100 lines without any additional text or explanations. - Population Density. - Land Density: Defined as the ratio of impervious surface area to the available land area.
4.4 Jane Jacobs’ urban vitality
Urban vitality theory is a classic urban theory put forward by Jacobs [36]. She pointed out that urban vitality often depends on four key elements: concentration (population density), functional diversity (architectural function mixture), contact opportunity (short block), and need for aged buildings (proportion of old buildings). These factors together constitute the foundation of urban vitality, among which population density and architectural function mixing are the core driving forces, while short blocks and old buildings enhance urban vitality by increasing contact opportunities and reducing operating costs. However, in recent years, there have been disputes about Jacobs’ views on high-rise buildings, and some studies believe that it has a positive impact on urban vitality [53].
Although Jacobs’ theory is widely accepted, empirical tests remain limited and inconsistent. The reasons for uncertainty are the robustness of the research design and the interpretability of the theory to cross-geographical areas. According to Jane Jacobs’ urban vitality Theory, there is a significant positive correlation between urban vitality and given factors. $P$ stands for Population Density, $B$ stands for Building Mix Index, $S$ stands for Short Block, $A$ stands for Aged Building, and $T$ stands for Tall Building.
$P=f(P,B,S,A,T,e)$ (10)
The understanding of Jane Jacobs’ urban vitality theory by the LLMs was tested. By inputting the appropriate prompt, the large model is used to generate urban samples, including land serial number, population density, building mixing degree, street intersection density, the proportion of old buildings and building height [53], and LLM is called as a resident for the second time to judge the livability of a block, so as to be the representative value of vitality, fit the theoretical equation, observe the goodness of fit and regression coefficient, and compare the differences of different large models. And compared with the numerical value of empirical research [54]. Each model repeats the independent experiment 10 times to generate data samples with averaged metrics.
Table 4: Prompt and specific description for urban vitality theory.
Prompt Description
Urban vitality theory – generation Generate attributes for 100 blocks in a city. Format the output as: Block Name1, Population Density1, Building Mix Index1, Short Block1, Aged Building1, Tall Building1 Block Name2, Population Density2, Building Mix Index2, Short Block2, Aged Building2, Tall Building2 Output exactly 100 lines without any additional text or explanations. - Population Density: People per square meter. - Building Mix Index: Entropy index based on building use. Return value 0 and 1. - Short Block: Total number of street intersections divided by the area of the block. - Aged Building: The rate of old buildings on each block. Return value 0 and 1. - Tall Building: The average building heights on each block.
Urban vitality theory – evaluation You are a resident of a city. Assign a livability score between 0 and 1 for each block, 0 represents the least livable and 1 represents the most livable. Evaluate each block based on the following attributes: - Block name. - Population Density: People per square meter. - Building Mix Index: Entropy index based on building use. - Short Block: Total number of street intersections divided by the area of the block. - Aged Building: The rate of old buildings on each block. - Tall Building: The average building heights on each block.
Competing Interests
All authors declare no financial or non-financial competing interests.
Fundings
This work was supported by the National Natural Science Foundation of China [Grant No. 62394331] and [Grant No. 62394335].
Data Availability
We have provided information on all publicly available data used in our analysis in the Methods.
Code Availability
The code to replicate the results in the paper will be available via Figshare at ”https://doi.org/10.6084/m9.figshare.28910084”.
Supplementary Information
Supplementary Note. Models Selection
For model selection, we evaluated top MEGA-Bench [21] models and conducted experiments to validate their real-world stability and applicability. While some models excelled in mathematical reasoning and code generation, they proved unsuitable for urban science data generation due to shortcomings in output completeness, stability, and controllability. For example:
Doubao-Pro-128k: The model produced incomplete outputs and failed to strictly follow the required prompt structure during experiments.
Hunyuan-Large-LongContext: The model produced incomplete outputs and exhibited high repetition across multiple runs, preventing diverse data generation. Additionally, it imposed usage frequency restrictions, impacting the reproducibility of batch experiments.
Qwen-Plus-Latest: The model produced incomplete outputs and failed to generate the required content in full, leading to missing information and reducing its suitability for rigorous data validation tasks.
ChatGPT, Claude, Gemini, DeepSeek, and ChatGLM demonstrated superior data generation completeness, theoretical alignment, and experimental reproducibility. Future research may further explore ways to optimize the application of other models to better meet the demands of urban science studies.
Supplementary Experiment 1. Comparison of Model Ability in the Data Distribution
As evidenced in Figure, GPT-4o and DeepSeek-V3 demonstrated superior alignment with real-world data among the five evaluated models. Under the JSD metric, these two models exhibited aggregate mean divergence values of 0.115 (R-G) and 0.116 (R-GC), representing reductions of $41.9\%$, $23.2\%$, and $1\%$ compared to Claude-3.5, Gemini-2.0, and ChatGLM-4, respectively.
Real data (R) comprises statistical records of population, GDP, and road network length across 685 Chinese municipal administrative regions. Each iteration involved random sampling of 100 cities, with this process repeated 100 times to ensure statistical robustness. Generated data (G) were generated by corresponding models, producing 100 city-level datasets per iteration over 100 cycles. The distinction between R-G origin and R-GC lies in the inclusion of the geographical cue ’China’ in the latter’s generation prompts. Pairwise Jensen-Shannon Divergence (JSD) analyses were conducted to quantify discrepancies among three datasets: real data (R), baseline generated data (G), and generated data with geographical prompt ’China’ (GC).
Supplementary Experiment 2. Human Height Data
We evaluated the ability of large language models (LLMs) to generate human height data. Given that height distributions in the real world follow well-documented statistical patterns, such as normal distributions with known mean and standard deviation values for different populations, this task serves as a baseline assessment of whether LLM-generated data aligns with real-world human height distributions. We conducted ten independent trials for each LLM to generate male and female height datasets. The generated data were then compared against the global average height distributions for males (mean = 178.4 cm, SD = 5.6 cm) and females (mean = 164.7 cm, SD = 4.7 cm), as reported by Silverman (2022). LLM-generated height distributions generally conformed to a normal pattern, though accuracy varied notably across models. Claude-3.5 consistently delivered the most realistic estimates, closely matching both the mean and standard deviation of true height data across genders. GPT-4o followed closely, offering stable performance with slightly conservative estimates. In contrast, DeepSeek-V3 exhibited the largest deviations, producing overly dispersed distributions that failed to capture real-world variance. Gender-based differences were also observed. For male height, GPT-4o achieved the best fit, with the highest R² (0.95) and near-identical distribution characteristics, while Claude-3.5 and ChatGLM-4 also performed well. For female height, Claude-3.5 was the most accurate overall, with the smallest deviation from the true mean and standard deviation, followed by ChatGLM-4, which demonstrated the best curve fit (R² = 0.95).
Supplementary Experiment 3. Individuals’ Wealth Data
We evaluated LLMs’ ability to generate individual-level wealth data and observed that the resulting distributions generally conformed to a Pareto pattern, capturing the right-skewed nature of wealth concentration. However, the estimated Pareto coefficients were consistently lower than the empirical benchmark of  1.5, as reported by Atkinson [atkinson2014], suggesting that LLMs tend to understate the extent of real-world wealth inequality. Despite this, the log-log linear fits between rank and wealth yielded relatively high R² values, indicating that LLMs can reproduce the structural form of the Pareto distribution. However, high R² does not imply accurate calibration of inequality levels, as the slope of the fitted line (i.e., the Pareto coefficient) remained biased toward more equal distributions. Among the tested models, Gemini-2.0 generated the most realistic distribution, producing both Pareto coefficients and tail behaviors closest to empirical values, thereby better reflecting the heavy-tailed nature of actual wealth distributions. Real-world wealth decile data used for comparison were obtained from the World Inequality Database [wid2025].
Figure S1: Comparative model capabilities based on R-G data difference.
Figure S2: Main LLMs tests: Human height distribution.
Figure S3: Main LLMs tests: Individuals’ wealth data distribution.
Figure S4: Sensitivity test of different levels of prompts.
Figure S5: Main LLMs tests of Distance decay theory: In terms of population density.
Figure S6: Main LLMs tests of Distance decay theory: In terms of urban land density.
Figure S7: Main LLMs tests of Zipf’s Law between cities.
Figure S8: Main LLMs tests of Urban Scaling Law in cities: infrastructure variables.
Figure S9: Main LLMs tests of Urban Scaling Law in cities: socio-economic variables.
Figure S10: GPT-4o with the best performance of urban scaling law.
Figure S11: Main LLMs tests of Jane Jacobs’ urban vitality theory. E = effective significant positive slope, P ¡ 0.05 and slope ¿ 0.

### v3

@@
-LLMs Capture Urban Science but Oversimplify Complexity
+GenAI Models Capture Urban Science but Oversimplify Complexity
 Abstract
-Large language models (LLMs) are increasingly used in scientific reasoning and data generation tasks, yet their alignment with empirical knowledge and practical urban data in urban science remains unclear.
-Here, this study introduces AI4US (Artificial Intelligence for Urban Science), a framework for systematically evaluating five leading LLMs (ChatGPT, DeepSeek, Claude, Gemini, and ChatGLM) by generating urban data and testing their fidelity against real data distributions and the empirical results of three foundational theories, including urban scaling laws, urban distance decay, and urban vitality.
-LLMs reproduce core theoretical patterns with notable accuracy (average $\mathrm{R}^{2}=0.804$ for scaling; $0.988$ for decay), and exhibit capacity to simulate less formalized constructs such as urban vitality.
-However, generated data exhibit crucial limitations: poor diversity compared to real data; systematic deviations in fitted theoretical parameters (e.g., underestimated scaling exponents and decay rates); and limited improvement from prompt engineering.
-These findings suggest that while LLMs offer a promising research paradigm for exploring urban science, they fall short of capturing real-world complexity, underscoring the need for improved alignment and evaluation.
+Generative artificial intelligence (GenAI) models are increasingly used for scientific data generation, yet their alignment with empirical knowledge in urban science remains unclear.
+Here, we introduce AI4US (Artificial Intelligence for Urban Science), a framework that systematically evaluates leading GenAI models by testing their fidelity in generating both symbolic and perceptual urban data.
+For the symbolic domain, we benchmark generated data against foundational urban theories concerning scale, space, and morphology.
+For the perceptual domain, we validate the models’ visual judgments against human benchmarks and, critically, leverage their generative control to conduct in causal experiments on urban perception.
+Our findings show that while GenAI models reproduce core theoretical patterns, the generated data exhibit crucial limitations: poor diversity, systematic parametric deviations, and improvement from prompt engineering.
+To address this, we introduce a post-hoc calibration procedure using optimal transport, which produces synthetic symbolic datasets with demonstrably higher fidelity.
 Machine Learning, ICML
 1 Introduction
-Urban science has long sought to decipher the complex behaviors of cities from early demographic surveys to computational modeling [1, 2, 3, 4].
-In recent decades, an explosion of urban data from remote sensing, sensor networks, and digital platforms has promised to revolutionize this endeavor [5, 6].
-Yet, a fundamental paradox remains.
-Despite this data abundance, theoretical innovation has stagnated [7, 8, 9].
-The field struggles to move beyond fragmented, context-dependent insights toward broadly generalizable urban theory.
-This impasse stems from a persistent methodological gap.
-The field lacks scalable and systematic experimental designs that allow for robust, cross-city replication and validation of scientific claims [8, 10, 11].
+Urban science has long sought to decipher the complex behaviors of cities from early demographic surveys to computational modeling [1, 2, 3, 4], which manifest across multiple scales (macro, meso, micro) and dimensions [5, 6].
+The endeavor to formulate urban theory from empirical observation has traditionally relied upon two principal data modalities: symbolic data, which abstracts urban phenomena into text and numerics, and perceptual data, which captures the direct sensory experience of the urban environment through modalities like visual imagery [7, 8, 9].
+Since the recent explosion of urban data [5, 6] from remote sensing, sensor networks, and digital platforms promised a revolution, struggling to move beyond fragmented, context-dependent insights toward broadly generalizable urban theory.
+However, theoretical innovation has stagnated [7, 8, 9].
+This impasse stems from a persistent methodological gap: the lack of scalable and systematic experimental designs that allow for robust, cross-city replication and validation of scientific claims [8, 10, 11].
 This theoretical stagnation is rooted in the inherent difficulty of conducting controlled experiments in urban settings.
@@
 These shortcomings are magnified by the profound transformations of contemporary cities, from autonomous mobility to platform economies, creating an urgent need for a new experimental paradigm capable of capturing the complex human-in-the-loop dynamics that define modern urban life [5, 19].
-Nowadays, advances in generative AI are spearheading a new paradigm of data-driven scientific discovery [20, 21, 22].
+Nowadays, advances in GenAI are spearheading a new paradigm of data-driven scientific discovery [20, 21, 22].
 In fields like biology and materials science, its success lies in predicting specific physical states from vast solution spaces, exemplified by AlphaFold’s protein structure prediction [23] and MatterGen’s discovery of novel materials [24].
 The challenge in urban science is fundamentally different.
-It is not to predict a discrete physical outcome [25], but to capture the broad statistical patterns that emerge from complex, adaptive urban systems, which are governed by context-dependent social behaviors rather than universal physical laws [26].
-This distinction positions LLMs as a uniquely compelling tool.
-Trained on diverse records of human behavior and expression, LLMs offer a powerful pathway to generate synthetic urban data with high fidelity, enabling scalable urban experimentation where formal modeling has reached its limits.
-In urban science, however, the potential of generative AI to function as a new form of virtual city laboratory that can produce experimental data [27, 28], analyze urban phenomena, and drive theoretical advancement remains largely unexplored [20, 29].
-Initial studies have shown promise, using LLMs to generate synthetic mobility data, simulate residents’ daily behaviors, or assess urban systems [22, 30].
+AI’s role in urban science is to capture the broad statistical patterns emerging from complex, adaptive urban systems governed by context-dependent social behaviors rather than universal physical laws [25].
+This distinction makes GenAI models particularly suitable.
+Their evolution from Large Language Models (LLMs), capable of generating and reasoning over symbolic data, to Multimodal Large Language Models (MLLMs) that also encompass perceptual data, unifies the two critical domains of urban information.
+Trained on diverse records of human behavior and expression, these models can thus generate high-fidelity synthetic data, enabling scalable urban experimentation where formal modeling has reached its limits.
+In urban science, however, the potential of GenAI to function as a new form of virtual city laboratory that can produce experimental data [26, 27], analyze urban phenomena, and drive theoretical advancement remains largely unexplored [20, 28].
+Initial studies have shown promise, using GenAI models to generate synthetic mobility data, simulate residents’ daily behaviors, or assess urban systems [22, 29].
 However, these pioneering efforts often remain fragmented.
-A significant research gap persists in systematically validating the fidelity of these worlds generated by AI against established urban theories [31].
-Furthermore, the inherent geographical and social biases within LLMs pose a critical challenge to their reliability as scientific instruments [32, 33].
-Therefore, a structured and rigorous framework is necessary to move beyond isolated applications and systematically evaluate the potential and pitfalls of LLMs as a foundation for the next generation of urban science.
+A significant research gap persists in systematically validating the fidelity of these worlds generated by GenAI against established urban theories [30].
+Furthermore, the inherent geographical and social biases within GenAI models pose a critical challenge to their reliability as scientific instruments [31, 32].
+Therefore, a structured and rigorous framework is necessary to move beyond isolated applications and systematically evaluate the potential and pitfalls of GenAI as a foundation for the next generation of urban science.
 To address these gaps, we introduce the AI4US (Artificial Intelligence for Urban Science) framework.
-World models are internal, generative representations that simulate real-world dynamics [33].
-This framework pioneers a methodology to evaluate LLMs as a first-generation proxy for urban world models.
-Our approach probes the implicit urban knowledge embedded within LLMs by testing their ability to reproduce three foundational urban theories: urban scaling laws, distance decay relationships, and Jane Jacobs’ urban vitality theory [29, 34, 35].
-These theories were chosen to create a comprehensive testbed, as they represent fundamental urban mechanisms across different scales and dimensions: urban scaling laws govern macro level regularities, distance decay explains meso scale spatial interactions, and urban vitality captures micro scale human environment dynamics [36, 37, 38, 39, 40, 41].
-Our findings reveal that while LLMs can replicate these theoretical patterns with notable fidelity, they also exhibit systematic limitations, underscoring that the path toward true theoretical discovery will require a new generation of world models capable of deeper causal reasoning.
+This framework establishes a ‘Generate-Evaluate-Calibrate’ workflow to systematically evaluate generative AI as a first-generation proxy by probing the implicit urban knowledge embedded within them.
+Our diagnostic approach evaluates the fidelity of AI-generated data across both symbolic and perceptual domains.
+For symbolic data, we test the models’ ability to reproduce foundational urban theories like urban scaling laws, distance decay, and urban vitality [28, 29, 30].
+For perceptual data, we assess their multimodal capacity to generate realistic urban scenes and replicate human-centric visual perception.
+Crucially, this generative control enables a form of in causal experimentation, allowing for systematic tests of how specific urban elements impact human perception.
+These theories were chosen to create a comprehensive testbed, as they represent fundamental urban mechanisms across different scales and dimensions [31, 32, 33, 34, 35, 36].
+Moving beyond diagnosis, this study introduces a novel post-hoc calibration procedure, demonstrating how a small sample of real data can be leveraged to rapidly assess these biases for a given urban dataset and subsequently correct the distribution of the larger generated dataset.
+This computationally efficient approach enhances the fidelity for downstream tasks, such as training predictive models or seeding simulations, without requiring costly modifications to the foundational GenAI models.
+Our findings reveal that the path toward the future will require advanced models [32].
 2 Results
-2.1 LLMs capture the fundamental forms of urban theories
-Figure 1: Urban scaling law reproduced on generated data with LLMs: generated versus empirical coefficients.
-“Replicate Tests” contains 100 independent duplicate data generation experiments for each LLM, and each experiment generates data of 100 cities, among which GPT-4o has the best performance (see Supplementary Experiment 6).
-The diagram shows the results of GPT-4o.
-Figure 2: Urban decay reproduced on generated data with LLMs: generated versus empirical statistical and spatial patterns.
-a, Concentric-ring land-density profiles for three representative cities on different continents across three decades (1990–2010), together with the corresponding “Mirage Cities” produced by GPT-4o when both city name and time period are included in the prompt.
-b, Sensitivity of the generated decay curves to prompts that specify different maximum radii and historical periods.
-c, Direct side-by-side comparison of Beijing’s empirical profile with that of its Mirage City generated by LLMs.
-Figure 3: Urban vitality revealed on generated data with LLMs.
-“Replicate Tests” contains 100 independent duplicate data generation experiments for each LLM, and each experiment generates data of 100 blocks.
-The diagram shows the results of GPT-4o.
-Figure 4: Data distribution divergence analysis.
-(a) Equal-width binning with relative frequency normalization was applied to convert raw sample counts into scale-invariant frequency distributions.
-Mean Absolute Error (MAE) of bin-level frequencies quantified the central tendency divergence between real data (R) and generated data (G), calculated as the sum of absolute differences across all 15 bins.
-Overlap Ratio (OR) measured quantile interval congruence between corresponding bins of R and G distributions.
-(b) Jensen-Shannon Divergence (JSD) comparisons among four datasets: real data (R), baseline generated data (G), China-prompted generated data (GC), and USA-prompted generated data (GA).
-The inclusion of geographical constraints, such as “China”, increased GPT-4o’s constraint-triggering rate by 24% (24/100 trials), e.g., “I can’t provide real-time data such as the current population”.
-Figure 5: AI4US, the proposed framework utilizing LLMs to advance urban science.
-The investigation reveals that LLMs possess a remarkable ability to recognize and reproduce the foundational forms of established urban theories across multiple scales, from macro-level city systems to micro-level neighborhood dynamics.
-At the macro-scale, when tasked with generating data for urban scaling laws, LLMs consistently produce outputs that adhere to the characteristic power-law relationship.
+2.1 GenAI models capture the fundamental statistical forms of urban theories
+The investigation reveals that GenAI models possess a strong ability to recognize and reproduce the foundational forms of established urban theories across multiple scales, from macro-level city systems to micro-level neighborhood dynamics.
+At the macro-scale, when tasked with generating data for urban scaling laws, GenAI models consistently produce outputs that adhere to the characteristic power-law relationship.
 The generated data exhibits a strong fit to the scaling law equation, achieving a high average coefficient of determination ($R^{2}=0.804$) (Fig.
@@
 Most models successfully replicated the expected super-linear relationship ($\beta>1$) for socioeconomic output and, in the case of GPT-4o, the sub-linear relationship for infrastructure.
+Figure 1: Urban scaling law reproduced on generated data with GenAI models: generated versus empirical coefficients.
+’Replicate Tests’ contains 100 independent duplicate data generation experiments for each GenAI model, and each experiment generates data of 100 cities, among which GPT-4o has best performance (See Supplementary Experiment 6).
+The diagram shows the results of GPT-4o.
 Moving to the meso-scale, the models display a non-trivial grasp of the concentric-zone structure posited by distance-decay theory.
-When generating land density surfaces, the outputs fit the inverse-S formulation with exceptional accuracy, showing a mean $R^{2}$ of 0.988 (Fig.
+When generating land density surfaces, the outputs fit the inverse-S formulation with high accuracy, showing a mean $R^{2}$ of 0.988 (Fig.
 2a).
-Furthermore, this capability extends beyond quantitative laws to less-formalized, qualitative theories.
-At the micro-scale, when confronted with urban vitality theory, LLMs demonstrate a clear alignment with its core tenets.
-By generating data for key indicators and then assessing their correlation with a livability score, the models produced relationships directionally consistent with Jane Jacobs’ theory (Fig.
-3a).
-Specifically, indicators such as ’Building Mix Index’ and ’Tall Building’ showed the highest consistency with the theory, while ’Building Mix Index’ and ’Population Density’ demonstrated the most significant positive impact on the generated livability scores.
-In summary, LLMs can consistently generate data that is structurally ”form-similar” to a diverse set of urban theories, whether they are mathematically explicit or qualitative, establishing a baseline of theoretical competence.
+Figure 2: Urban decay reproduced on generated data with GenAI models: generated versus empirical statistical and spatial patterns.
+a, Concentric-ring land-density profiles for three representative cities on different continents across three decades (1990-2010), together with the corresponding ’Mirage cities’ produced by GPT-4o when both city name and time period are included in the prompt.
+b, Sensitivity of the generated decay curves to prompts that specify different maximum radii and historical periods.
+c, Direct side-by-side comparison of Beijing’s empirical profile with that of its mirage city generated by GenAI models.
+Furthermore, this capability extends to descriptive theories that lack a single mathematical form, such as Jane Jacobs’ theory of urban vitality.
+The test of this theory reveals a distinct capability of GenAI, where the model is prompted to act as a ”resident” to simulate a ”social consensus” on what makes a place livable.
+The model’s outputs revealed nuanced and theoretically consistent insights, identifying ’Building Mix Index’ and ’Tall Building’ as having the highest alignment with the theory, while ’Aged Building’ showed the greatest deviation (Fig.
+3).
+This demonstrates that an GenAI model can process the complex, multi-faceted information embedded in qualitative urban theories to generate data that is structurally ”form-similar” and directionally consistent with their core tenets.
+More importantly, it shows a capacity to translate these qualitative concepts into structured, testable hypotheses regarding the relative importance of different urban elements, establishing a baseline of theoretical competence.
+Figure 3: Urban vitality revealed on generated data with GenAI models.
+’Replicate Tests’ contains 100 independent duplicate data generation experiments for the GenAI models, and each experiment generates data of 100 blocks.
+The diagram shows the results of GPT-4o.
+2.2 A new paradigm for qualitative theory exploration in visual urban space
+This research demonstrates that GenAI models can function as a new paradigm for exploring and validating theories in visual urban space.
+The fidelity of the generated images was first confirmed in a blind expert validation test.
+AI-generated images received a mean realism score of 7.83 (SD = 1.6), which was not statistically different from the 7.68 (SD = 1.8) scored by real photographs (t(198) = 0.65, p = 0.52).
+This high fidelity was complemented by substantial diversity, achieved via a structured-prompting strategy that increased the average CLIP semantic distance to 0.3310.
+This represents a 35.8% improvement over a baseline single-prompt approach ($D_{CLIP}=0.2438$) and surpasses the diversity of a 1,000-image real-world benchmark ($D_{CLIP}=0.2905$), enabling the creation of semantically rich datasets.
+Having established the quality of the generated data, we validated the AI’s ability to emulate human perception and identify its visual drivers.
+The AI’s perceptual judgments showed substantial agreement with human choices from the Place Pulse dataset across all six evaluated categories, with Kappa values consistently above 0.2 and agreement rates approaching 70% (Fig.
+4a).
+Using this validated AI perception, we then quantified the influence of visual elements on these attributes through a multiple linear regression model.
+The analysis revealed several significant relationships (Fig.
+4b).
+For instance, the presence of a ’person’ was a strong positive predictor for ’lively’ ($\beta=0.223,p<0.001$), while features such as ’wall’ negatively impacted the perception of ’beautiful’ ($\beta=-0.247,p<0.001$).
+To further validate these patterns, we compared our results to the benchmark findings of Place Pulse [37], which shows that ’tree’ and ’grass’ are strong positive predictors for ’beautiful’, while ’sky’ is a significant negative predictor for ’safe’.
+Finally, the ”generate-modify-evaluate” paradigm was used to test the AI’s utility for causal inference by quantifying the impact of thematic interventions (Fig.
+4c).
+The addition of Natural Elements produced the most pronounced effect, significantly increasing the perceived ’beautiful’ score by an average of 2.5 points ($\Delta$ Score) and having the strongest negative impact on the ’depressing’ score.
+Conversely, adding Traffic Elements was the most effective intervention for increasing the ’lively’ score, demonstrating the framework’s ability to test the perceptual outcomes of specific design strategies.
+Figure 4: Potential of GenAI models in qualitative theory exploration in visual urban space.
+a, AI perception aligns with human judgment.
+A confusion matrix of pairwise choices (left) and per-category agreement rates with Cohen’s Kappa scores (right).
+b, Identifying influential visual elements.
+Standardized beta coefficients ($\beta$) from regression models reveal the key visual features that are positively or negatively correlated with each of the six perceptual scores.
+c, Causal inference via thematic interventions.
+The average change in perception scores ($\Delta$ Score) quantifies the causal impact of programmatically adding thematic element categories to scenes: Natural Elements (trees, grass, sky, water, bushes), Traffic Elements (cars, trucks, bridges, roads), and Built Elements (buildings, fences, walls).
+Error bars represent 95% confidence intervals.
 2.3 Mirage cities systematically oversimplify real-world complexity
-While LLMs can replicate theoretical forms, a deeper analysis reveals a significant gap between their generated ”Mirage cities” and the complexity of real urban systems.
+While GenAI models can replicate theoretical forms based on statistical data, a deeper analysis reveals a significant gap between their generated ”Mirage cities” and the complexity of real urban systems.
 This discrepancy manifests as a series of systematic, non-random deviations.
 A primary finding is the lack of diversity in the generated data compared to empirical data.
-The analysis of data distributions reveals that LLM-generated data is significantly more homogeneous.
+The analysis of data distributions reveals that generated data is significantly more homogeneous.
 Without specific geographic prompts, the generated data’s numerical scale diverges substantially from real data, resulting in a low average Overlap Ratio (OR) of just 28.6% per bin (Fig.
-4a).
+5a).
 More critically, the Jensen-Shannon Divergence (JSD) between real-world data samples (R-R) is, on average, 122.54% higher than the divergence within generated samples (G-G) (Fig.
-4b).
+5b).
 This indicates that real data exhibits substantially greater internal variation.
-This inherent ”over-smoothing” tendency is a foundational reason why LLM-generated data, while theoretically clean, lacks the real-world noise and heterogeneity essential for robust scientific analysis.
+This inherent ”over-smoothing” tendency is a foundational reason why generated data, while theoretically clean, lacks the real-world noise and heterogeneity essential for robust scientific analysis.
 This lack of diversity contributes to specific parametric and structural distortions when fitting theories.
@@
 The distributional difference in these generated decay curves is 149% smaller than that of empirical data, further underscoring their idealized nature.
+This pattern of oversimplification extends to the visual domain.
+The AI’s perceptual judgment, while directionally correct, showed weaknesses in abstract cases, achieving lower consistency on subjective concepts like ’boring’ ($\kappa=0.155$) and ’lively’ ($\kappa=0.157$).
+Most notably, the AI exhibited a strong bias against ambiguity, agreeing with human ’equal’ judgments in only 2.1% of cases and instead forcing a definitive choice (Fig.
+4a).
+Figure 5: Data distribution divergence analysis.
+a, Equal-width binning with relative frequency normalization was applied to convert raw sample counts into scale-invariant frequency distributions.
+Mean Absolute Error (MAE) of bin-level frequencies quantified the central tendency divergence between real data (R) and generated data (G), calculated as the sum of absolute differences across all 15 bins.
+Overlap Ratio (OR) measured quantile interval congruence between corresponding bins of R and G distributions.
+b, Jensen-Shannon Divergence (JSD) comparisons among four datasets: real data (R), baseline generated data (G), China-prompted generated data (GC), and USA-prompted generated data (GA).
+The inclusion of geographical constraints, such as ’China’, increased GPT-4o’s constraint-triggering rate by 24% (24/100 trials) (e.g., ’I can’t provide real-time data such as the current population’).
 2.4 The potential for improving AI-generated data
+A post-hoc calibration procedure based on Optimal Transport (OT) is shown to be the most effective intervention in our comparative study for enhancing data fidelity (Supplementary Experiment 8).
+Our proposed OT method demonstrates consistent improvements over the raw generative output, reducing the Jensen-Shannon Divergence (JSD) for the GDP distribution from approximately 0.6 to 0.35 and increasing the Overlapping Ratio (OR) for infrastructure by a factor of five (Supplementary Experiment 8a).
+More critically, the calibration rectifies the systematic parametric biases found in the raw generative output.
+For instance, it corrects the Zipf’s Law exponent from a biased value of approximately -0.5 to a value that aligns with the theoretical -1.0, and similarly adjusts the scaling exponents for infrastructure and GDP to match their empirical benchmarks (Supplementary Experiment 8b).
+The method’s effectiveness is further validated by an ablation study, which shows that applying OT to uninformed noise (Noise + OT) yields only limited improvements, confirming the synergistic importance of the GenAI’s sophisticated generative prior.
+Furthermore, the OT consistently outperforms the traditional statistical synthesizer (KDE), positioning it as a robust method for generating high-fidelity urban data.
+2.5 The impact of different prompts and models
 Given the observed deviations, this study investigated whether these gaps could be bridged through different forms of intervention.
@@
 Prompting with specific geographic contexts demonstrates a clear effect on data distribution; when prompted with ”China,” the generated data showed a greater alignment with real data from Chinese cities, increasing the Overlap Ratio by 38.56% and decreasing the Mean Absolute Error by 52.14% (Fig.
-4a, 4b).
+5a, 5b).
 Similarly, at the meso-scale, prompting with parameters like city size or historical period directly influences the resulting decay curves (Fig.
 2c).
-These results confirm that the model’s output is not static but sensitive to contextual cues, highlighting a clear potential for iterative refinement.
-Furthermore, performance varies noticeably across the different LLMs evaluated.
+These results demonstrate that the model’s output is sensitive to contextual cues, highlighting a clear potential for iterative refinement.
+Furthermore, performance varies noticeably across the different GenAI models evaluated.
 In the urban scaling law test, GPT-4o demonstrated the closest alignment with empirical results, while Claude 3.5 was most accurate in reproducing Zipf’s law.
@@
 These variations suggest that factors intrinsic to the models, such as architecture, training data, and fine-tuning strategies, significantly influence the characteristics and fidelity of the generated urban data.
-2.4 A new paradigm for qualitative theory exploration
-The test of urban vitality, a theory that is descriptive and lacks a single mathematical form, reveals a distinct capability of LLMs, shifting their role from a ”physical simulator” to a ”socio-cultural simulator.” By prompting an LLM to act as a ”resident” evaluating a neighborhood, the task is not to solve a physical equation, but to have the model tap into its vast training data of human text and discourse to simulate a ”social consensus” on what makes a place livable.
-The model’s outputs revealed nuanced and theoretically consistent insights (Fig.
-3).
-For instance, it identified ’Building Mix Index’ and ’Tall Building’ as having the highest consistency with vitality theory, while ’Aged Building’ showed the greatest deviation.
-In terms of impact, it judged ’Building Mix Index’ and ’Population Density’ as having the most significant positive effect on livability scores (Fig.
-3b).
-This demonstrates that LLMs can serve as a powerful ”hypothesis generation engine.” They can process the complex, multi-faceted information embedded in qualitative theories and produce structured, testable hypotheses about the relative importance of different urban elements.
-This capability to ”operationalize” qualitative knowledge marks a significant new potential for AI in advancing urban science.
 3 Discussion
-This study reveals a fundamental duality in the capabilities of current Large Language Models within urban science.
-They excel at reproducing the abstract, mathematical forms of established theories but falter in capturing the statistical complexity and fidelity of real world urban systems.
-The systematic evaluation across different scales demonstrates that while LLMs can generate data that is structurally ”form similar” to urban theories, this apparent competence masks significant deficiencies in diversity and parametric accuracy.
-These findings position LLMs not as high fidelity physical simulators, but as powerful engines for exploring the socio-cultural patterns embedded in human language, a role that offers both profound opportunities and critical limitations for scientific inquiry.
-The initial finding that LLMs consistently generate ”mirage cities” with low diversity and systematically distorted parameters (e.g., underestimated scaling exponents) points to a core limitation.
+This study reveals a fundamental duality in the capabilities of current GenAI models within urban science.
+They excel at reproducing the abstract, mathematical forms of established statistical theories but falter in capturing the complexity of real-world urban systems.
+This pattern of oversimplification extends directly to the visual domain.
+The systematic evaluation across different scales demonstrates that while GenAI models can generate data that is structurally ”form-similar” to urban theories, this apparent competence masks significant deficiencies.
+In the statistical realm, this manifests as a lack of diversity and systematically distorted parameters (e.g., underestimated scaling exponents).
+In the visual realm, it appears as a bias against perceptual ambiguity (agreeing with human ’equal’ judgments in only 2.1% of cases) and a failure to capture the full variance of global urban environments.
+The initial finding that GenAI models consistently generate ”mirage cities” with low diversity and systematically distorted parameters (e.g., underestimated scaling exponents) points to a core limitation.
 This phenomenon, where models produce archetypal representations of cities, a tendency sometimes described as ”group flattening” [38], aligns with observations of geographical and social bias in other domains [31, 32].
-This interpretation suggests that these distortions are not random errors but are symptomatic of the LLMs’ nature as statistical pattern matchers rather than causal world models.
+This interpretation suggests that these distortions are symptomatic of the GenAI’s nature as statistical pattern matchers rather than causal world models.
 Lacking a deep, generative understanding of the mechanisms that produce urban phenomena, they instead reproduce the most dominant statistical patterns from their training data, smoothing over the granular heterogeneity and outliers that are often crucial drivers of real world dynamics.
-This tendency, also noted in other applications of LLMs for generating synthetic populations [31], poses a significant risk for empirical research that relies on capturing the full spectrum of urban complexity.
-The investigation into data fidelity reveals that LLMs are not static instruments but are malleable tools sensitive to user interaction and model specific characteristics.
-Our finding that prompt engineering can partially mitigate deviations by steering the model toward a specific context underscores the importance of the human AI interface in scientific applications [30, 31].
-The quest is not for a single, perfect prompt, but for a methodology of iterative refinement where the prompt itself is a key parameter in the experimental design [42].
-Furthermore, the observed performance differences across the five evaluated LLMs highlight that model architecture and training strategies are critical variables.
-This suggests that future improvements in data fidelity will depend on both more sophisticated interaction protocols and the rapid advancements in the models themselves [20].
-The limitations of LLMs also unveil a novel scientific opportunity.
-Their tendency to generate theoretically ”clean” data, stripped of real world noise, makes them powerful instruments for theoretical exploration [29, 34, 35, 31, 33].
+This tendency, also noted in other applications of GenAI models for generating synthetic populations [30], poses a significant risk for empirical research that relies on capturing the full spectrum of urban complexity.
+The investigation into data fidelity reveals that GenAI models are malleable tools sensitive to user interaction and model specific characteristics.
+Our finding that prompt engineering can partially mitigate deviations by steering the model toward a specific context underscores the importance of the human AI interface in scientific applications [29, 30, 39].
+The observed performance differences across the five evaluated GenAI models suggests that future improvements in data fidelity will depend on both more sophisticated interaction protocols and the rapid advancements in the models themselves [20].
+Beyond these interactive refinements, our successful implementation of a post-hoc calibration procedure using Optimal Transport offers a crucial insight: the raw output of an GenAI model should be seen as a structurally rich yet uncalibrated starting point.
+This hybrid approach effectively overcomes the limitations of relying solely on the model’s internal world representation, demonstrating a practical path toward creating high-fidelity synthetic datasets.
+A distinct capability of GenAI models is revealed when they are tasked with navigating less formalized, qualitative theories like urban vitality.
+Here, the GenAI is prompted to simulate a ”social consensus” by synthesizing the vast amount of information about cities embedded in its training data, from academic texts to public discourse.
+The result is the generation of a plausible, structured set of relationships that can provide urban scientists with novel, testable hypotheses.
+This aligns with the growing recognition of GenAI models as tools for urban sensing and knowledge synthesis.
+In this role, the GenAI acts as a powerful catalyst for human led scientific inquiry, offering a new method to ”operationalize” qualitative knowledge into a quantitative, empirical research program.
+The GenAI models also unveil a novel scientific opportunity.
+Their tendency to generate data by guidance of prompts with controlling variables in experiments, makes them powerful instruments for theoretical exploration [28, 29, 30, 31, 32].
 Urban scientists can now generate vast, fine grained datasets under specific theoretical assumptions, creating virtual laboratories to explore a theory’s logical consequences or test its sensitivity to certain parameters.
@@
 This approach directly addresses the challenge of how AI can generate new knowledge for urban studies, by providing a scalable method to rigorously interrogate theoretical constructs in a controlled environment.
-A distinct capability of LLMs is revealed when they are tasked with navigating less formalized, qualitative theories like urban vitality.
-Here, the LLM is not solving a physical equation but is prompted to simulate a ”social consensus” by synthesizing the vast amount of information about cities embedded in its training data, from academic texts to public discourse.
-The result is not a ”discovery” of a new law, but the generation of a plausible, structured set of relationships that can provide urban scientists with novel, testable hypotheses.
-This aligns with the growing recognition of LLMs as tools for urban sensing and knowledge synthesis.
-In this role, the LLM acts as a powerful catalyst for human led scientific inquiry, offering a new method to ”operationalize” qualitative knowledge into a quantitative, empirical research program.
+However, while our structured prompting strategy enhanced the diversity of the generated outputs, a fundamental limitation persists: the current methodology is incapable of synthesizing a truly de novo virtual city and is instead constrained to generating data corresponding to random, extant locations on Earth.
+While GenAI models can be prompted to generate data for hypothetical or counterfactual scenarios, interpreting these outputs as scientifically valid predictions is challenging.
+For example, correcting known biases in generated data to test unknown scenarios where data cannot be obtained in the real world.
+The primary obstacle is the ”black box” nature of the models.
+We cannot audit their internal reasoning process.
+Furthermore, the systematic biases we measure in one context, such as under baseline conditions, are not guaranteed to be stable when the model is prompted with a novel, out-of-distribution scenario.
+Therefore, these counterfactual generations are most valuable when used to produce structured and testable hypotheses that guide subsequent empirical research.
 While this study provides a foundational framework, its scope defines a clear agenda for future research.
-Continuous benchmarking is essential, as the quantitative performance of LLMs will evolve with new architectures [20].
+Continuous benchmarking is essential, as the quantitative performance of GenAI models will evolve with new architectures [20].
 More importantly, our findings underscore a critical distinction.
-The systematic biases and lack of causal depth suggest that current LLMs, while powerful, are not true world models.
+The systematic biases and lack of causal depth suggest that current GenAI models, while powerful, are not true world models.
 They are adept at learning the ”what” from data but struggle with the ”why”.
 The path toward transformative breakthroughs in urban science, particularly in areas like causal inference and policy simulation [40], will likely require a new generation of hybrid models.
-Future work should focus on integrating the rich, contextual knowledge of LLMs with the causal integrity of traditional urban simulation frameworks, a synergy that could unlock the next frontier of a robust and predictive urban science.
+These findings position GenAI models as powerful engines for synthesizing urban scientific data and exploring the socio-cultural patterns for scientific inquiry.
+Future work should focus on integrating the rich, contextual knowledge of GenAI models with the causal integrity of traditional urban simulation frameworks, a synergy that could enable new avenues for a robust and predictive urban science.
 4 Method
-The intelligence of large language models essentially derives from their ability to compress world knowledge acquired during training [44, 45].
-To systematically evaluate the capabilities of LLMs in urban science, this study employs a framework grounded in three foundational urban theories (Fig.
-5).
-The selection of these theories is deliberate, designed to test the LLMs’ generative fidelity across the distinct yet interconnected scales at which urban phenomena are understood: the macro-level system of cities, the meso-level internal structure of a single city, and the micro-level dynamics of a neighborhood.
+The intelligence of GenAI models essentially derives from their ability to compress world knowledge acquired during training [41, 42].
+To systematically evaluate the capabilities of GenAI models in urban science, this study employs a framework that tests their generative fidelity against two core domains of urban theory: (1) statistical laws based on symbolic data (different dimensions with scale, space, and morphology), and (2) cognitive laws based on perceptual data (such as street-view images in urban scenes).
+Figure 6: AI4US, the proposed framework utilizing GenAI models to advance urban science.
+4.1 Urban theories based on generating symbolic data
+For the first domain, we selected three foundational theories designed to test the GenAI’s abilities across the distinct yet interconnected scales at which urban phenomena are understood: the selection of these theories is deliberate, designed to test the GenAI’s generative fidelity across the distinct yet interconnected scales at which urban phenomena are understood: the macro-level system of cities, the meso-level internal structure of a single city, and the micro-level dynamics of a neighborhood (Supplementary Note 2).
 Substantial evidence suggests that, similar to regularities observed in natural complex systems, urban phenomena exhibit patterns and order across these multiple dimensions of size, space, and morphology [36, 43, 44, 45].
-The chosen theories are canonical representations of these scaled phenomena (Table 1).
+The chosen theories are canonical representations of these scaled phenomena (Table S1).
 Urban scaling laws capture nonlinear regularities at the macro-scale, describing how infrastructure and socioeconomic activity grow with city size.
@@
 Together, these theories exemplify foundational mechanisms in urban systems.
-They provide a comprehensive and representative testbed for evaluating the core capabilities of LLMs: from reproducing mathematically explicit laws to simulating the socio-cultural knowledge embedded in qualitative theory.
-Table 1: A Multi-Scale Framework for Evaluating LLMs in Urban Science.
-Scale Theory Significance in Urban Science
-Macro Urban scaling law Explains universal, size-based regularities of cities as
-(City system) complex systems, central to understanding urban growth
-and efficiency.
-Meso Urban distance decay Represents the foundational role of spatial friction and
-(Internal city structure) accessibility in shaping internal urban structure and
-organization.
-Micro Urban vitality Provides a human-centric perspective linking the micro-design
-(Block) of the built environment to social and economic outcomes;
-a cornerstone of urban design.
+They provide a comprehensive and representative testbed for evaluating the core capabilities of GenAI models: from reproducing mathematically explicit laws to simulating the socio-cultural knowledge embedded in qualitative theory.
+4.1.1 Urban Scaling Law
+As a foundational concept in complex systems, urban scaling laws describe the power-law relationship between a city’s size and its various metrics, serving as functional models for understanding cities as complex systems.
+These laws are broadly categorized into inter-urban processes (interactions between cities) [46], such as Zipf’s Law, and intra-urban processes (dynamics within a city) that explore how human interactions and spatial geometry contribute to scaling.
+Zipf’s Law posits an inverse relationship between a city’s population and its rank.
+$\log_{10}(\text{Population})=\log_{10}a+\beta\log_{10}(\text{Rank})$ (1)
+The intra-urban scaling law relates socioeconomic or infrastructural metrics $y$ to city population.
+$\log_{10}(y)=\log_{10}a+\beta\log_{10}(\text{Population})$ (2)
+$\log_{10}(a)$ represents the intercept, while the slope $\beta$ is the critical scaling exponent.
+Empirical evidence [38] suggests two distinct patterns for intra-urban scaling: socioeconomic activities (e.g., GDP) exhibit super-linear scaling ($\beta>1$), implying increasing returns to scale, while infrastructure metrics show sub-linear scaling ($\beta<1$), indicating greater efficiency in larger cities.
+To test these laws, GenAI models were prompted to generate a dataset of 100 cities with attributes for population, infrastructure, and GDP.
+4.1.2 Urban Distance Decay
+This theory addresses the meso-scale structure of cities, positing that the intensity of interaction between two locations diminishes with increasing distance, a concept closely related to Tobler’s First Law of Geography [47, 48].
+This typically results in a pronounced ”center-periphery” pattern in urban spatial structures, where attributes like population density and land value decrease from the city core.
+While several quantitative models exist, our study focused on the inverse S-function model for characterizing the spatial decay of urban land density $f(r)$ as a function of distance ($r$) from the city center [49, 50, 51].
+$f(r)=\frac{1-c}{1+e^{a\left(\frac{2r}{D}-1\right)}}+c$ (3)
+The parameters define the curve’s shape: $c$ represents the minimum land density in the urban periphery (the lower asymptote); $a$ is the growth rate, controlling the steepness of the density decay; and $D$ is a scaling parameter related to the city’s spatial extent, where the inflection point of the curve occurs at the distance $r=D/2$.
+To assess the models’ grasp of this concept, GenAI models were prompted to generate land density values for 100 concentric urban zones.
+4.1.3 Jane Jacobs’ Urban Vitality
+At the micro-scale, we tested the GenAI’s ability to operationalize Jane Jacobs’ classic theory of urban vitality [35].
+The theory posits that vitality depends on four key elements: concentration, functional diversity, contact opportunity (short blocks), and aged buildings, where population density and building mix are considered core driving forces [52].
+Despite its wide acceptance, empirical tests of the theory remain limited and inconsistent [huang2024b], and debates persist around certain elements, such as the role of tall buildings.
+$P=f(P,B,S,A,T,e)$ (4)
+Here, vitality is a function of Population Density ($P$), Building Mix Index ($B$), Short Block ($S$), Aged Building ($A$), and Tall Building ($T$).
+We employed a two-step process: first, GenAI models were prompted to generate data for 100 urban blocks across these five key attributes; second, the models were prompted to act as a ”resident” and assign a ”livability score” to each block as a proxy for vitality, allowing us to test whether the models’ inferred relationships were consistent with Jacobs’ core tenets.
+4.2 Urban perception through generating perceptual data
+For the second domain, we evaluated the GenAI’s multimodal capabilities in urban visual perception.
+We focused on human-scale street view imagery (SVI), a mature benchmark representing the daily urban life.
+The evaluation used a ”Generate-Evaluate-Calibrate” paradigm grounded in the six perceptual indicators from the MIT Place Pulse project [37, 53, 54, 55, 56, 57] (Table S2).
+Our methodology involved a three-phase process to validate the AI as a visual data generator, a perceptual judge, and a tool for causal inference.
+Figure 7: A computational framework for advancing urban perception through perceptual data. Edit text and images through Nano Banana.
+First, to validate the AI’s data generation capability, the generated human-scale streetscapes were assessed for realism and diversity.
+Realism was evaluated by human experts, who rated 100 images from AI generation and 100 images from Place Pulse on a 1-10 scale (Supplementary Experiment 9).
+The model’s fidelity was then quantified by comparing the score distributions against real photographs using an independent samples t-test.
+Dataset diversity was ensured via a structured-prompting strategy.
+The effectiveness of this strategy was then quantitatively validated by using CLIP embeddings to compare the average semantic distance $D_{CLIP}$ across three distinct image sets: a real-world street view benchmark from Place Pulse, a baseline set generated from simple prompts, and our proposed structured-prompt set.
+$D_{CLIP}=\frac{2}{N(N-1)}\sum_{i=1}^{N}\sum_{j=i+1}^{N}\sqrt{1-\frac{e_{i}\cdot e_{j}}{\|e_{i}\|\|e_{j}\|}}$ (5)
+Second, to validate the AI’s perceptual judgment, its pairwise choices on original Place Pulse images were compared against the source dataset’s human choices using Kappa.
+$K=\frac{P_{0}-P_{e}}{1-P_{e}}$ (6)
+Furthermore, to assess if the AI could identify influential visual factors in alignment with human patterns [57], we fit a multiple linear regression model correlating the AI’s perception scores ($Y_{p}$) with the pixel proportions of 150 visual elements ($X_{k}$) from semantic segmentation.
+$Y_{p}=\beta_{p,0}+\sum_{k=1}^{K}\beta_{p,k}X_{k}+\varepsilon_{p}$ (7)
+Third, to validate the AI’s utility for causal inference, we performed thematic interventions on baseline images to create ”before-and-after” pairs.
+These interventions were designed around three interpretable categories: Natural, Traffic, and Built Elements.
+The causal effect of each category was then quantified by calculating the average change in perception scores ($\Delta S_{p,c}$) between the variant and baseline images for each perception type.
+$\Delta S_{p,c}=\frac{1}{N_{c}}\sum_{i=1}^{N_{c}}(S_{p,i}^{\text{variant}}-S_{p,i}^{\text{baseline}})$ (8)
+Where $\Delta S_{p,c}$ is the average change in score for perception $p$ under category $C$, $S_{p,i}^{\text{variant}}$ and $S_{p,i}^{\text{baseline}}$ are the scores for a single image pair $i$, and $N_{c}$ is the total number of image pairs in that category.
 4.3 Comparison of data distribution
 First, this study directly compares the data distributions of generated and real-world samples.
-Since LLM-generated values without geographic prompts often differ from real samples by orders of magnitude, this study applies equal-width binning and relative frequency normalization to eliminate differences in data scale, converting the raw sample counts within each bin into frequencies.
+Since generated values without geographic prompts often differ from real samples by orders of magnitude, this study applies equal-width binning and relative frequency normalization to eliminate differences in data scale, converting the raw sample counts within each bin into frequencies.
 The Mean Absolute Error (MAE) of bin-level frequencies is then used to measure the average difference in relative frequency between real and generated data across bins, serving as an indicator of how well the generated data approximates the overall distribution trend of the real data.
@@
 $\text{Overlap Ratio}=\frac{\sum_{i=1}^{N}\text{Intersection}(Q_{R,i},Q_{G,i})}{\sum_{i=1}^{N}\text{Union}(Q_{R,i},Q_{G,i})}$ (10)
-Secondly, based on the urban data generated by LLMs and the real data of Chinese cities, this study measures the pairing difference between the two data and three discrete distributions by Jensen-Shannon Divergence (JSD).
+Secondly, based on the urban data generated by GenAI models and the real data of Chinese cities, this study measures the pairing difference between the two data and three discrete distributions by Jensen-Shannon Divergence (JSD).
 Include ’distance between real sample and real sample data distribution’ (R-R), ’distance between real sample and generated sample data distribution’ (R-G), ’Distance between generated sample and generated sample data distribution’(G-G).
@@
 $\text{Distance}_{R-G}$ $=\frac{1}{n\times m}\sum_{i=1}^{n}\sum_{j=1}^{m}\text{JSD}(R_{i},G_{j})$ (11)
-$\text{JSD}(R,G)$ $=\frac{1}{2}\text{KL}(R\|\frac{1}{2}(R+G))+\frac{1}{2}\text{KL}(G\|\frac{1}{2}(R+G))$ (4)
+$\text{JSD}(R,G)$ $=\frac{1}{2}\text{KL}(R\|M)+\frac{1}{2}\text{KL}(G\|M)$ (12)
+$\text{where }M$ $=\frac{1}{2}(R+G)$
 $\text{KL}(R\|M)$ $=\sum_{x}R(x)\log\frac{R(x)}{M(x)}$ (13)
-4.2 Urban scaling law
+4.4 Post-hoc calibration procedure for data distribution
+To address the systematic bias and lack of diversity observed in raw generated data, we introduce a post-hoc calibration procedure (detailed explanations refer to Supplementary Note 3).
+This challenge of aligning a generated (’source’) distribution with an empirical (’target’) distribution is a fundamental problem in machine learning, analogous to domain adaptation [58, 59] or batch effect correction [60].
+While prior work in generated data synthesis has focused on prompt engineering or simple rule-based adjustments [26, 30, 31], these methods may not correct for systemic distributional shifts.
+We therefore employ an approach grounded in Optimal Transport (OT) theory [58, 61] for its ability to transform an entire dataset while preserving its core structure.
+The procedure learns an optimal transport map $T$ that transports samples from the preprocessed generated distribution ($P_{G}$) to the preprocessed real data distribution ($P_{real}$).
+This map is then applied to the full generated dataset ($x_{i}$) to produce a calibrated set ($D_{\text{calib, scaled}}$):
+$D_{\text{calib, scaled}}=\{T(x_{i})|x_{i}\in D_{G,\text{scaled}}\}$ (14)
+Here, $x_{i}$ is the set of generated data points, and T is the learned mapping function.
+The distance between distributions is quantified using the Wasserstein distance ($W_{1}$), which represents the minimum ”cost” required to transform one distribution into another:
+$W_{1}(P_{real},P_{G})=\inf_{\gamma\in\Pi(P_{real},P_{G})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|x-y\|d\gamma(x,y)$ (15)
+Where $\Pi(P_{real},P_{G})$ is the set of all possible joint distributions of the source $P_{G}$ and target ($P_{real}$) distributions.
+The efficacy of this OT approach was validated against several baselines, including raw GenAI output, random noise, and data synthesized via KDE.
+This comprehensive correction is achieved because the OT procedure learns a mapping that transforms the entire joint distribution of the multi-dimensional data.
+By adjusting the data structure holistically, rather than treating each variable independently, it inherently corrects the inter-variable correlations that define the scaling exponents.
+Competing Interests
+All authors declare no financial or non-financial competing interests.
+Data Availability
+We have provided information on all publicly available data used in our analysis in the Methods. Please refer to Table S3 for all prompts in this study.
+Code Availability
+The code to replicate the results in the paper will be available via Figshare.
+Supplementary Information
+Supplementary Note 1. Models selection
+For model selection, we evaluated top MEGA-Bench [21] models and conducted experiments to validate their real-world stability and applicability.
+While some models excelled in mathematical reasoning and code generation, they proved unsuitable for urban science data generation due to shortcomings in output completeness, stability, and controllability.
+For example:
+  • Doubao-Pro-128k: The model produced incomplete outputs and failed to strictly follow the required prompt structure during experiments.
+  • Hunyuan-Large-LongContext: The model produced incomplete outputs and exhibited high repetition across multiple runs, preventing diverse data generation.
+Additionally, it imposed usage frequency restrictions, impacting the reproducibility of batch experiments.
+  • Qwen-Plus-Latest: The model produced incomplete outputs and failed to generate the required content in full, leading to missing information and reducing its suitability for rigorous data validation tasks.
+ChatGPT, Claude, Gemini, DeepSeek, and ChatGLM demonstrated superior data generation completeness, theoretical alignment, and experimental reproducibility.
+Future research may further explore ways to optimize the application of other models to better meet the demands of urban science studies.
+Supplementary Note 2. Detailed explanations for selected theories
+Urban scaling law
 Scaling laws are a fundamental concept in complex systems, describing how certain quantities change with system size, typically following a power-law relationship.
@@
 For intra-urban processes, empirical evidence suggests two distinct scaling laws: (1) variables related to socioeconomic activity (e.g., GDP, patents, HIV cases) exhibit super-linear scaling ($\beta>1$), meaning that as a city’s population increases, these socioeconomic indicators grow at a faster rate; and (2) variables related to infrastructure (e.g., the number of charging stations or gas stations) exhibit sub-linear scaling ($\beta<1$), indicating that as the population increases, these infrastructure-related variables grow at a slower rate.
-For Zipf’s Law between cities, the relationship typically follows a power-law function where $\beta\approx-1$, meaning that city size (e.g., population) and city rank follow an inverse relationship.
-In this equation, $\log_{10}(a)$ represents the intercept, corresponding to the constant $a$ in the power-law function, while the slope $\beta$ determines the exponent of the power-law relationship.
-Here, $y$ represents population size, and $x$ denotes city rank.
-The model incorporates several statistical parameters: $n$ refers to the number of independent trials fitting the theoretical equation, $g$ represents the number of generated data samples per trial, and $m$ denotes the parameters of the theoretical equation (such as $a$ and $\beta$ in Zipf’s Law between cities).
-The standard deviation of $m$ across $n$ trials reflects variability in parameter estimation, while the mean value of $m$ provides an overall estimate.
-A narrower confidence interval (CI) for $\beta$ suggests a more precise estimate.
-Additionally, the coefficient of variation (CV) for Zipf’s Law parameters is computed as the geometric mean of the CV for $a$ and $\beta$, assuming that these parameters are independent and without considering their correlations.
-$\log_{10}(\text{Population})=\log_{10}a+\beta\log_{10}(\text{Rank})$ (6)
-$R^{2}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y}_{i})^{2}}{\sum_{i=1}^{n}(y_{i}-\bar{y})^{2}}$ (7)
-The urban scaling law in cities describes how the infrastructure and socio-economic factors in a city change with the growth of the city scale.
-It is the scaling law of urban infrastructure and urban social factors derived by Bettencourt in 2013 based on the Human Interaction model.
-$Y$ represents the infrastructure or socio-economic elements in the city.
-$\text{CV}_{\text{Urban}}$ represents the coefficient of variation of the parameters of the urban scaling law in cities under $n$ independent experiments.
-$\log_{10}(y)=\log_{10}a+\beta\log_{10}(\text{Population})$ (8)
-The understanding of Zipf’s Law and urban scaling law by the LLMs was tested for inter-city and intra-city models, respectively.
-By inputting the appropriate prompt, the large model is used to generate city samples, including city serial number, population, road length, and GDP, and the theoretical equations of inter-city model and urban internal model are fitted respectively, and the goodness of fit and regression coefficient are observed, and the differences between different large models are compared.
-And compare it with the value of empirical research [38].
-Each model repeats the independent experiment 10 times to generate data samples with averaged metrics.
-Table 2: Prompt and specific description for urban scaling law.
-Prompt Description
-Urban scaling law Generate the dataset containing 100 cities in a country.
-Format the output as: CityName1, Population1, Infrastructure volume1, GDP1 CityName2, Population2, Infrastructure volume2, GDP2 Output exactly 100 lines without any additional text or explanations.
-- City Name - Population - Infrastructure volume: Total road miles - GDP: all Gross Domestic Product of the city in one year
-4.3 Urban distance decay
+Urban distance decay
 The term distance decay is used to describe the diminishing intensity of interaction between two locations as the distance between them increases.
@@
 For urban land use density, the inverse S-function model has also been proposed to characterize its spatial decay [51].
-For urban land density, the inverse S-function model describes a decaying distribution pattern that follows an inverse S-shaped curve as one moves outward from the city center.
-In this model, $L$ represents the saturation density parameter, indicating the maximum possible land density (i.e., the asymptote of the curve).
-$k$ denotes the growth rate parameter (1/km), which reflects how rapidly the density transitions from low values to the saturation level; the larger the value of $k$, the steeper and more abrupt the transition.
-$x_{0}$ is the center point parameter (km), representing the location at which the land density reaches half of its maximum value — essentially, the inflection points of the curve.
-$\mathrm{CV}_{\mathrm{DLand}}$ represents the coefficient of variation of the related parameters of land density attenuation in the distance decay theory under $n$ independent experiments.
-$f(r)=\frac{1-c}{1+e^{a\left(\frac{2r}{D}-1\right)}}+c$ (9)
-The extent to which current LLMs understand the spatial distribution of population density and land density according to the distance decay law was assessed.
-By providing appropriate prompts, LLMs simulated a mirage city with 100 concentric zones and generated samples that included zone identifiers, population density, and land density attributes for each zone.
-Then the study fitted theoretical equations for both population and land density to the generated data, evaluating the goodness of fit and regression coefficients, and compared these values with those found in empirical studies [52].
-Additionally, the study examined the differences across various LLMs.
-Each model conducted 100 independent trials to generate data samples, and finally reported the averaged results across these repetitions.
-Table 3: Prompt and specific description for urban distance decay theory.
-Prompt Description
-Urban distance decay theory Generate attributes for 100 concentric city circles arranged in order in a city.
-From the city center (Circle 1) to the outermost ring (Circle 100).
-Format the output as: Circle 1, Population Density1, Land Density 1 Circle 2, Population Density2, Land Density 2 Output exactly 100 lines without any additional text or explanations.
-- Population Density.
-- Land Density: Defined as the ratio of impervious surface area to the available land area.
-4.4 Jane Jacobs’ urban vitality
+Jane Jacobs’ urban vitality
 Urban vitality theory is a classic urban theory put forward by Jacobs [35].
@@
 The reasons for uncertainty are the robustness of the research design and the interpretability of the theory to cross-geographical areas.
-According to Jane Jacobs’ urban vitality Theory, there is a significant positive correlation between urban vitality and given factors.
-$P$ stands for Population Density, $B$ stands for Building Mix Index, $S$ stands for Short Block, $A$ stands for Aged Building, and $T$ stands for Tall Building.
-$P=f(P,B,S,A,T,e)$ (10)
-The understanding of Jane Jacobs’ urban vitality theory by the LLMs was tested.
-By inputting the appropriate prompt, the large model is used to generate urban samples, including land serial number, population density, building mixing degree, street intersection density, the proportion of old buildings and building height [53], and LLM is called as a resident for the second time to judge the livability of a block, so as to be the representative value of vitality, fit the theoretical equation, observe the goodness of fit and regression coefficient, and compare the differences of different large models.
-And compared with the numerical value of empirical research [54].
-Each model repeats the independent experiment 10 times to generate data samples with averaged metrics.
-Table 4: Prompt and specific description for urban vitality theory.
+Urban perception through images
+The AI4US framework extends its evaluation beyond statistical regularities to the domain of urban visual perception.
+Urban science is concerned not only with emergent patterns but also with the human-centric experience of the built environment, which is profoundly shaped by its visual characteristics.
+Understanding the link between physical appearance and human psychological responses, such as feelings of safety or aesthetic pleasure, is fundamental to urban design and planning.
+The study of urban perception has evolved significantly over the past half-century.
+It originated from foundational, qualitative approaches, such as Kevin Lynch’s work on the urban image [62], which relied on manual surveys and expert analysis to decode how people form mental maps of their cities.
+The subsequent integration of photography and GIS allowed researchers to begin spatializing these subjective perceptions.
+A paradigm shift occurred with the rise of computational methods and large-scale crowdsourcing, exemplified by the MIT Place Pulse project [37, 53, 54, 55, 56].
+This initiative collected millions of human judgments on street view imagery (SVI), creating a global dataset that links images to perceptual scores.
+This enabled the next stage: using machine learning to automate perception analysis at an unprecedented scale, and leveraging Explainable AI (XAI) to quantify which visual elements correlate with specific perceptions [57].
+The current era of generative multimodal models marks a new frontier, allowing for end-to-end experiments where scenes can be generated, programmatically modified, and re-evaluated to test causal hypotheses about perception.
+Supplementary Note 3. Post-hoc calibration procedure for data distribution
+To overcome the systematic bias and lack of diversity observed when using GenAI to synthesize data, it is necessary to calibrate the larger dataset generated by GenAI by utilizing a small portion of real empirical data.
+This challenge of aligning a generated (’source’) distribution with an empirical (’target’) distribution is a fundamental problem in machine learning [58].
+It is conceptually analogous to domain adaptation in computer vision, where synthetic data is transformed to match real-world image characteristics [59], or to batch effect correction in bioinformatics, where data from different experiments are normalized to a common distribution [60].
+In these contexts, the goal is to learn a mapping that corrects for distributional shifts while preserving the core data structure.
+Within the specific context of GenAI-based data synthesis, prior research has largely focused on enhancing output fidelity through sophisticated prompt engineering or agent-based architectures [30, 31].
+Post-processing, when applied, has typically involved direct, rule-based adjustments.
+These include enforcing strict output formats and value ranges to ensure data validity at the point of generation [30], or subsequently grounding the GenAI’s conceptual outputs (e.g., a ”trip to the office”) in physical space by mapping them to specific geographic coordinates based on real-world constraints like a city’s road network [26].
+While these methods improve the plausibility of individual data points, they may not correct for systematic biases in the overall data distribution.
+To address this need, we selected an approach grounded in Optimal Transport (OT) [58, 61] theory for its distinct advantages over simpler alternatives.
+Direct probability density estimation methods (e.g., Kernel Density Estimation) often struggle with the curse of dimensionality in multi-attribute urban data.
+Simpler filtering methods, such as classifier-based selection, can discard a significant portion of the generated samples and may fail to preserve the latent structure of the data.
+In contrast, OT offers a more powerful solution by transforming the entire generated dataset rather than merely selecting from it, providing a robust method for distributional alignment that is well-suited to this task.
+Table S1: A Multi-Scale Framework for Evaluating GenAI models in Urban Science.
+Scale Theory Significance in Urban Science
+Macro (City system) Urban scaling law Explains universal, size-based regularities of cities as complex systems, central to understanding urban growth and efficiency.
+Meso (Internal city structure) Urban distance decay Represents the foundational role of spatial friction and accessibility in shaping internal urban structure and organization.
+Micro (Block) Urban vitality Provides a human-centric perspective linking the micro-design of the built environment to social and economic outcomes; a cornerstone of urban design.
+Table S2: Benchmark for Urban Perception Indicators based on Visual Elements.
+Summary of regression coefficients ($\beta$) for all unique top-10 visual objects influencing six urban perceptions.
+The significance of each coefficient is indicated by stars.
+Adapted from the analysis in Zhang et al.
+(2018) [57].
+The top 10 objects that positively/negatively contributed to each of the 6 perception types are shown.
+Beta Coefficient (*p¡0.1,**p¡0.05,***p¡0.01).
+Visual Object Safe Lively Beautiful Wealthy Depressing Boring
+Bridge - -0.0144*** -0.0120** - 0.0178*** 0.0076
+Building -0.1163*** -0.0625*** -0.0689*** - 0.0483*** -
+Car 0.0600*** 0.0887*** - 0.0295*** -0.0243*** -0.0329***
+Earth - - - -0.0174*** - -
+Fence -0.0327*** - -0.0201*** - 0.0198*** 0.0144***
+Field - -0.0543*** - -0.0248*** - 0.0130**
+Floor - - -0.0172*** -0.0219*** - -
+Grass 0.0711*** - 0.0465*** 0.0272*** -0.0204*** -
+Hill - - - - - 0.0105*
+House - - - - -0.0123** -
+Path 0.0154*** - 0.0126*** 0.0115** - -
+Pole - - - - - -0.008
+Road 0.0395*** 0.0228*** - 0.0122* - -
+Signboard - - - - - 0.0173***
+Sidewalk 0.0821*** 0.0596*** - - -0.0350*** -0.0320***
+Sky -0.2068*** -0.2389*** -0.1179*** -0.1142*** 0.1063*** 0.0969***
+Tree -0.0392*** -0.1055*** 0.0980*** 0.0244*** -0.0610*** -
+Truck - - -0.0125*** - - -
+Wall -0.0764*** -0.0828*** -0.0458*** -0.0264*** 0.0389*** 0.0382***
+Water - - 0.0070* - - -
+Table S3: Prompts and specific descriptions in this study.
 Prompt Description
+Urban scaling law Generate the dataset containing 100 cities in a country.
+Format the output as: CityName1, Population1, Infrastructure volume1, GDP1 CityName2, Population2, Infrastructure volume2, GDP2 Output exactly 100 lines without any additional text or explanations.
+- Infrastructure volume: Total road miles - GDP: all Gross Domestic Product of the city in one year
+Urban distance decay theory Generate attributes for 100 concentric city circles arranged in order in a city.
+From the city center (Circle 1) to the outermost ring (Circle 100).
+Format the output as: Circle 1, Population Density1, Land Density 1 Circle 2, Population Density2, Land Density 2 Output exactly 100 lines without any additional text or explanations.
+- Population Density.
+- Land Density: Defined as the ratio of impervious surface area to the available land area.
 Urban vitality theory – generation Generate attributes for 100 blocks in a city.
-Format the output as: Block Name1, Population Density1, Building Mix Index1, Short Block1, Aged Building1, Tall Building1 Block Name2, Population Density2, Building Mix Index2, Short Block2, Aged Building2, Tall Building2 Output exactly 100 lines without any additional text or explanations.
+Format the output as: Block Name1, Population Density1, Building Mix Index1, Short Block1, Aged Building1, Tall Building1 Output exactly 100 lines without any additional text or explanations.
 - Population Density: People per square meter.
 - Building Mix Index: Entropy index based on building use.
-Return value 0 and 1.
+Value 0 to 1.
 - Short Block: Total number of street intersections divided by the area of the block.
 - Aged Building: The rate of old buildings on each block.
-Return value 0 and 1.
+Value 0 to 1.
 - Tall Building: The average building heights on each block.
@@
 - Building Mix Index: Entropy index based on building use.
-- Short Block: Total number of street intersections divided by the area of the block.
+- Short Block: Total number of street intersections of the block.
 - Aged Building: The rate of old buildings on each block.
 - Tall Building: The average building heights on each block.
-Competing Interests
-All authors declare no financial or non-financial competing interests.
-Fundings
-This work was supported by the National Natural Science Foundation of China [Grant No. 62394331] and [Grant No. 62394335].
-Data Availability
-We have provided information on all publicly available data used in our analysis in the Methods.
-Code Availability
-The code to replicate the results in the paper will be available via Figshare at ”https://doi.org/10.6084/m9.figshare.28910084”.
-Supplementary Information
-Supplementary Note. Models Selection
-For model selection, we evaluated top MEGA-Bench [21] models and conducted experiments to validate their real-world stability and applicability.
-While some models excelled in mathematical reasoning and code generation, they proved unsuitable for urban science data generation due to shortcomings in output completeness, stability, and controllability.
-For example:
-Doubao-Pro-128k: The model produced incomplete outputs and failed to strictly follow the required prompt structure during experiments.
-Hunyuan-Large-LongContext: The model produced incomplete outputs and exhibited high repetition across multiple runs, preventing diverse data generation.
-Additionally, it imposed usage frequency restrictions, impacting the reproducibility of batch experiments.
-Qwen-Plus-Latest: The model produced incomplete outputs and failed to generate the required content in full, leading to missing information and reducing its suitability for rigorous data validation tasks.
-ChatGPT, Claude, Gemini, DeepSeek, and ChatGLM demonstrated superior data generation completeness, theoretical alignment, and experimental reproducibility.
-Future research may further explore ways to optimize the application of other models to better meet the demands of urban science studies.
+Streetscape Generation (Phase 1 Base Template) ”A photorealistic, eye-level street view of {scene_description} from a random place on earth.
+The screen should be completely filled with the generated image, without white borders.
+Try to approach the real scene captured by the car (A specialized vehicle equipped with professional camera equipment for capturing street view images) as closely as possible, without any retouching or beautification.
+The camera perspective is a forward-facing view, looking straight down the road.
+It should be a head-up view without any perspective distortion.”
+Causal Intervention (Phase 3 Base Template) Realistically edit the image to {intervention_description}. Keep everything else in the image the same.
 Supplementary Experiment 1. Comparison of Model Ability in the Data Distribution
-As evidenced in Figure, GPT-4o and DeepSeek-V3 demonstrated superior alignment with real-world data among the five evaluated models.
-Under the JSD metric, these two models exhibited aggregate mean divergence values of 0.115 (R-G) and 0.116 (R-GC), representing reductions of $41.9\%$, $23.2\%$, and $1\%$ compared to Claude-3.5, Gemini-2.0, and ChatGLM-4, respectively.
+As evidenced in Figure S1, GPT-4o and DeepSeek-V3 demonstrated superior alignment with real-world data among the five evaluated models.
+Under the JSD metric, these two models exhibited aggregate mean divergence values of 0.115 (R-G) and 0.116 (R-GC), representing reductions of 41.9%, 23.2%, and 1% compared to Claude-3.5, Gemini-2.0, and ChatGLM-4, respectively.
 Real data (R) comprises statistical records of population, GDP, and road network length across 685 Chinese municipal administrative regions.
@@
 Pairwise Jensen-Shannon Divergence (JSD) analyses were conducted to quantify discrepancies among three datasets: real data (R), baseline generated data (G), and generated data with geographical prompt ’China’ (GC).
+Figure S1: Comparative model capabilities based on R-G data difference.
 Supplementary Experiment 2. Human Height Data
-We evaluated the ability of large language models (LLMs) to generate human height data.
-Given that height distributions in the real world follow well-documented statistical patterns, such as normal distributions with known mean and standard deviation values for different populations, this task serves as a baseline assessment of whether LLM-generated data aligns with real-world human height distributions.
-We conducted ten independent trials for each LLM to generate male and female height datasets.
+In this supplementary experiment, we evaluated the ability of GenAI models to generate human height data.
+Given that height distributions in the real world follow well-documented statistical patterns, such as normal distributions with known mean and standard deviation values for different populations, this task serves as a baseline assessment of whether generated data aligns with real-world human height distributions.
+We conducted ten independent trials for each GenAI model to generate male and female height datasets.
 The generated data were then compared against the global average height distributions for males (mean = 178.4 cm, SD = 5.6 cm) and females (mean = 164.7 cm, SD = 4.7 cm), as reported by Silverman (2022).
-LLM-generated height distributions generally conformed to a normal pattern, though accuracy varied notably across models.
+Generated height distributions generally conformed to a normal pattern, though accuracy varied notably across models.
 Claude-3.5 consistently delivered the most realistic estimates, closely matching both the mean and standard deviation of true height data across genders.
@@
 For female height, Claude-3.5 was the most accurate overall, with the smallest deviation from the true mean and standard deviation, followed by ChatGLM-4, which demonstrated the best curve fit (R² = 0.95).
+Figure S2: Main GenAI models tests: Human height distribution.
 Supplementary Experiment 3. Individuals’ Wealth Data
-We evaluated LLMs’ ability to generate individual-level wealth data and observed that the resulting distributions generally conformed to a Pareto pattern, capturing the right-skewed nature of wealth concentration.
-However, the estimated Pareto coefficients were consistently lower than the empirical benchmark of  1.5, as reported by Atkinson [atkinson2014], suggesting that LLMs tend to understate the extent of real-world wealth inequality.
-Despite this, the log-log linear fits between rank and wealth yielded relatively high R² values, indicating that LLMs can reproduce the structural form of the Pareto distribution.
+We evaluated GenAI’s ability to generate individual-level wealth data and observed that the resulting distributions generally conformed to a Pareto pattern, capturing the right-skewed nature of wealth concentration.
+However, the estimated Pareto coefficients were consistently lower than the empirical benchmark of  1.5, as reported by Atkinson and Piketty, suggesting that GenAI models tend to understate the extent of real-world wealth inequality.
+Despite this, the log-log linear fits between rank and wealth yielded relatively high R² values, indicating that GenAI models can reproduce the structural form of the Pareto distribution.
 However, high R² does not imply accurate calibration of inequality levels, as the slope of the fitted line (i.e., the Pareto coefficient) remained biased toward more equal distributions.
 Among the tested models, Gemini-2.0 generated the most realistic distribution, producing both Pareto coefficients and tail behaviors closest to empirical values, thereby better reflecting the heavy-tailed nature of actual wealth distributions.
-Real-world wealth decile data used for comparison were obtained from the World Inequality Database [wid2025].
-Figure S1: Comparative model capabilities based on R-G data difference.
-Figure S2: Main LLMs tests: Human height distribution.
-Figure S3: Main LLMs tests: Individuals’ wealth data distribution.
+Real-world wealth decile data used for comparison were obtained from the World Inequality Database.
+Figure S3: Main GenAI models tests: Individuals’ wealth data distribution.
 Figure S4: Sensitivity test of different levels of prompts.
-Figure S5: Main LLMs tests of Distance decay theory: In terms of population density.
-Figure S6: Main LLMs tests of Distance decay theory: In terms of urban land density.
-Figure S7: Main LLMs tests of Zipf’s Law between cities.
-Figure S8: Main LLMs tests of Urban Scaling Law in cities: infrastructure variables.
-Figure S9: Main LLMs tests of Urban Scaling Law in cities: socio-economic variables.
+Figure S5: Main GenAI models tests of Distance decay theory: In terms of population density.
+Figure S6: Main GenAI models tests of Distance decay theory: In terms of urban land density.
+Figure S7: Main GenAI models tests of Zipf’s Law between cities.
+Figure S8: Main GenAI models tests of Urban Scaling Law in cities: infrastructure variables.
+Figure S9: Main GenAI models tests of Urban Scaling Law in cities: socio-economic variables.
 Figure S10: GPT-4o with the best performance of urban scaling law.
-Figure S11: Main LLMs tests of Jane Jacobs’ urban vitality theory. E = effective significant positive slope, P ¡ 0.05 and slope ¿ 0.
+Figure S11: Main GenAI models tests of Jane Jacobs’ urban vitality theory. E = effective significant positive slope, P ¡ 0.05 and slope ¿ 0.
+Figure S12: Enhancing the effectiveness of GenAI as a data synthesizer based on optimal transport.
+The top panels display the mean relative frequency distributions, while the bottom panels provide a quantitative evaluation using three key performance metrics.
+All results are aggregated from 100 independent experimental runs.
+In each run, a large synthetic dataset (1,000 data points) is generated by each method, calibrated against a new random sample of 100 real cities.
+Figure S13: Urban streetscape image realism evaluation.